### v2

# ADPO: Anchored Direct Preference Optimization A Unified Framework from Pairwise to Listwise Preferences

Abstract
Direct Preference Optimization (DPO) is an efficient alternative to reinforcement learning from human feedback (RLHF), yet it typically assumes hard binary labels and pairwise comparisons. Such assumptions can be brittle under noisy or distribution-shifted supervision. We present Anchored Direct Preference Optimization (ADPO), which (i) incorporates soft preference probabilities, (ii) aligns policy updates through reference anchoring that induces an implicit trust region, and (iii) extends to listwise learning via Plackett–Luce modeling. In controlled synthetic setups covering 12 scenarios (4 noise types $\times$ 3 severities) and 3 model scales, ADPO exhibits relative improvements ranging from 12% to 79% over a standard DPO baseline (10-seed means; 95% CIs in the Appendix). Hard labels tend to fare better under severe noise, whereas soft labels yield better calibration under distribution shift; listwise variants achieve the highest WinMass (expected probability mass on the ground-truth best item) in 9/12 scenarios. Larger models amplify ADPO’s benefits (0.718 vs. 0.416 at hidden=256), suggesting that anchoring acts as an effective trust-region regularizer. We release code and configurations to facilitate reproducibility.
1 Introduction
Background. Preference optimization has emerged as the dominant paradigm for aligning large language models (LLMs) with human values. Traditional RLHF [2, 3] learns an explicit reward model from preference comparisons, then optimizes the policy using PPO [4]. Direct Preference Optimization (DPO) [1] simplifies this pipeline by directly optimizing policy log-ratios to match preference probabilities, eliminating the reward modeling stage.
Problem. Standard DPO [1] makes two restrictive assumptions: (i) preferences are hard binary labels ($y_{ij}\in\{0,1\}$), ignoring uncertainty; (ii) comparisons are strictly pairwise, limiting expressiveness. While DPO includes a reference model $\pi_{\text{ref}}$ for implicit KL regularization via the log-ratio $\log\frac{\pi_{\theta}}{\pi_{\text{ref}}}$, its pairwise structure anchors only the difference $\Delta_{\theta,\text{ref}}=(\log\pi_{\theta}-\log\pi_{\text{ref}})_{y_{i}}-(\log\pi_{\theta}-\log\pi_{\text{ref}})_{y_{j}}$. This makes DPO sensitive to initialization and prone to overfitting noisy labels—when training data contains noise due to annotator disagreement, adversarial examples, or distribution shift, these limitations cause gradient drift and brittleness.
Key question. Does the choice between soft and hard labels matter? Previous work assumes soft labels provide noise robustness through confidence weighting, but lacks controlled experiments isolating this factor. Similarly, the benefit of reference anchoring has not been systematically quantified under varying noise regimes and model scales.
Contributions. This work (a) unifies pairwise and listwise preference learning with anchored log-odds matching, recovering DPO/BT/PL as special cases (Proposition 3.2); (b) provides a controlled, multi-scenario study disentangling anchoring and label softness, with ablations on temperatures, model scale, and reference initialization; and (c) offers practical guidance for method selection under distinct noise regimes. While our findings are derived from synthetic contextual bandit setups, the observed trends—especially the robustness gains from anchoring—should be informative for RLHF-style pipelines.
2 Related Work
Preference optimization. More generally, the Preference Optimization (PO) framework [17] formulates an entropy-regularized objective aligning policy probabilities with qualitative preference signals, which inspires our entropy-regularized interpretation of anchoring (Section 3), though our work focuses on preference optimization for RLHF-style alignment rather than combinatorial optimization. DPO [1] reparameterizes the reward model through policy-reference log-ratios, enabling direct optimization. Extensions include identity-free preferences (IPO [5]), contrastive objectives (CPO [6]), and iterative refinement [7]. Our work provides a unified framework encompassing these variants.
Listwise preference learning. Plackett–Luce models [8, 9] enable listwise preferences through recursive top-1 selections. Recent work applies PL to LLM alignment [10, 11, 12]. We extend DPO to listwise settings through reference anchoring.
Robust learning under noise. Prior work addresses noise through majority voting [14], uncertainty quantification [15], and robust reward learning [16]. Our approach encodes uncertainty in soft probabilities and provides anchoring for groupwise shift invariance.
Entropy-regularized grounding and its link to ADPO.
Under the maximum-entropy RL objective
$\max_{\pi}\ \mathbb{E}_{x\sim\mathcal{D}}\Big[\mathbb{E}_{\tau\sim\pi(\cdot|x)}r(x,\tau)\ +\ \alpha\,\mathsf{H}\big(\pi(\cdot|x)\big)\Big],$
the optimal policy admits the Boltzmann form [18, 19]
$\pi^{*}(\tau|x)\;=\;\frac{1}{Z(x)}\exp\!\Big(\tfrac{1}{\alpha}\,r(x,\tau)\Big),$ (1)
and the reward can be reparameterized by the (optimal) policy log-probability
$r(x,\tau)\;=\;\alpha\,\log\pi^{*}(\tau|x)\;+\;\alpha\,\log Z(x),$ (2)
so any relative quantity (pairwise/listwise) cancels $Z(x)$. Eqs. (1)–(2) yield a direct bridge between reward differences and policy log-odds: Bradley–Terry and Plackett–Luce targets arise by mapping $\Delta r$ to a preference probability via a sigmoid/softmax, while ADPO matches the student’s anchored log-odds $(s-s^{\mathrm{ref}})/\tau$ to these soft targets. A full derivation of (1) as the minimizer of $\mathbb{E}_{x}\,\mathrm{KL}\big(\pi(\cdot|x)\,\|\,\pi^{*}(\cdot|x)\big)$ and the log-partition cancellation in pairwise/listwise forms can be found in [PO2025] (see their Eqs. (3)–(6), (20)–(27)).
Lemma 2.1 (Groupwise shift invariance).
If $r^{\prime}(x,\tau)=r(x,\tau)-h(x)$ for any function $h$ independent of $\tau$, then $\pi^{*}$ in (1) is unchanged; hence all pairwise/listwise probabilities (BT/PL) are identical, and anchored ADPO—which operates on $(s-s^{\mathrm{ref}})$—inherits groupwise shift invariance. See [PO2025], Prop. 3.1 and App. D.2.
3 ADPO: Unified Formulation
3.1 Pairwise Soft-DPO
For a candidate pair $(i,j)$ with soft teacher preference $q_{ij}\in(0,1)$ that $i$ is preferred, the ADPO pairwise loss is:
$\ell_{ij}^{\text{Soft-DPO}}=\log\left(1+e^{\beta(\Delta_{\theta}-\Delta_{\text{ref}})}\right)-q_{ij}\,\beta(\Delta_{\theta}-\Delta_{\text{ref}}),$ (3)
where $\beta>0$ is student temperature, $\Delta_{\theta}=s_{i}-s_{j}$, and $\Delta_{\text{ref}}=s_{i}^{\text{ref}}-s_{j}^{\text{ref}}$ with $s_{i}=\log\pi_{\theta}(\tau_{i}|x)$.
Key properties:
  • Bayes-optimal matching: Gradient vanishes when $\sigma(\beta(\Delta_{\theta}-\Delta_{\text{ref}}))=q_{ij}$.
  • Soft weighting: Uncertain pairs ($q_{ij}\approx 0.5$) contribute less gradient.
  • Reference anchoring: Relative updates $\Delta_{\theta}-\Delta_{\text{ref}}$ are invariant to groupwise shifts.
Remark 3.1 (Scale non-identifiability).
Temperatures $\beta$ and $\beta_{r}$ only appear through ratio $\beta_{r}/\beta$ in the optimum. We set $\beta=\beta_{r}=1.0$ by default.
Proposition 3.2 (Special cases).
Soft-DPO recovers: (i) standard DPO when $q_{ij}\in\{0,1\}$ and $\Delta_{\text{ref}}=0$; (ii) non-anchored Bradley–Terry when $\Delta_{\text{ref}}=0$; (iii) reward-based soft preferences when $q_{ij}=\sigma(\beta_{r}(R_{i}-R_{j}))$.
Proof of (i).
For $q_{ij}=1$: $\ell_{ij}=\log(1+e^{\beta\Delta_{\theta}})-\beta\Delta_{\theta}=-\log\sigma(\beta\Delta_{\theta})$. For $q_{ij}=0$: $\ell_{ij}=\log(1+e^{\beta\Delta_{\theta}})=-\log\sigma(-\beta\Delta_{\theta})$. This matches standard DPO exactly. ∎
3.2 Listwise Soft-DPO
For group $S_{x}=\{\tau_{1},\ldots,\tau_{P}\}$, the ADPO listwise loss is:
$\mathcal{L}_{\text{group}}^{\text{ref}}=\mathbb{E}_{x,S_{x}}\left[-\sum_{i\in S_{x}}q(i|S_{x})\log\tilde{p}_{\theta}(i|S_{x})\right],$ (4)
where
$\tilde{p}_{\theta}(i|S_{x})=\frac{\exp\left((s_{i}-s_{i}^{\text{ref}})/\tau\right)}{\sum_{j\in S_{x}}\exp\left((s_{j}-s_{j}^{\text{ref}})/\tau\right)},$ (5)
and teacher target $q(i|S_{x})\propto\exp(\hat{R}_{i}/\beta_{r})$ uses transformed rewards $\hat{R}_{i}$ via: (a) raw: $\hat{R}_{i}=\tilde{R}_{i}$; (b) rank-based Gaussian transform; or (c) KDE-CDF-Logit: $\hat{R}_{i}=\mathrm{logit}(\hat{F}(\tilde{R}_{i}))$ where $\hat{F}$ is the KDE-estimated CDF.
Equivalent expansion. The anchored listwise cross-entropy (4) can be equivalently written as:
$\sum_{i\in S_{x}}q(i|S_{x})\log\tilde{p}_{\theta}(i|S_{x})=-\frac{1}{\tau}\sum_{i\in S_{x}}q(i|S_{x})\big(s_{i}-s_{i}^{\mathrm{ref}}\big)+\log\!\sum_{j\in S_{x}}\exp\!\left(\frac{s_{j}-s_{j}^{\mathrm{ref}}}{\tau}\right).$ (6)
This decomposition makes explicit that the anchored listwise cross-entropy consists of a linear matching term $\langle q,(s-s^{\mathrm{ref}})/\tau\rangle$ and a log-sum-exp normalization. It directly reveals the connection to the convex conjugate form of the softmax, underlying the implicit trust-region regularization discussed below. Since $\mathrm{KL}(q\|\tilde{p}_{\theta})=\mathcal{L}_{\text{group}}^{\text{ref}}+H(q)$ and $H(q)$ is parameter-independent, minimizing (6) is equivalent to minimizing the KL divergence from the teacher distribution $q$ to the student’s anchored distribution $\tilde{p}_{\theta}$.
Lemma 3.3 (Implicit trust region; quadratic form at $p=q$).
Let $u_{i}=(s_{i}-s_{i}^{\mathrm{ref}})/\tau$ and $\tilde{p}_{\theta}(i;u)=\frac{e^{u_{i}}}{\sum_{j}e^{u_{j}}}$. Let $u^{\star}$ satisfy $\tilde{p}_{\theta}(\cdot;u^{\star})=q(\cdot)$ and define the $q$-centered logits $\delta_{i}=u_{i}-\sum_{j}q_{j}u_{j}$. Then, as $u\to u^{\star}$,
$\mathrm{KL}\!\big(q\,\|\,\tilde{p}_{\theta}(\cdot;u)\big)\;=\;\tfrac{1}{2}\,\delta^{\top}\!\big(\mathrm{Diag}(q)-qq^{\top}\big)\,\delta\;+\;o(\|\delta\|^{2})\;=\;\frac{1}{2\tau^{2}}\,\mathrm{Var}_{q}\!\big[s-s^{\mathrm{ref}}\big]\;+\;o(\cdot),$
where the variance is taken with respect to $q$ and the quadratic form uses the softmax Fisher metric $\mathrm{Diag}(q)-qq^{\top}$ (which is invariant to adding constants to all logits).
Sketch.
Write $\mathcal{L}(u)=A(u)-\langle q,u\rangle$ with $A(u)=\log\sum_{j}e^{u_{j}}$. Then $\nabla A(u)=\tilde{p}_{\theta}(\cdot;u)$ and $\nabla^{2}A(u)=\mathrm{Diag}(\tilde{p}_{\theta})-\tilde{p}_{\theta}\tilde{p}_{\theta}^{\top}$. A second-order Taylor expansion of $\mathcal{L}$ at $u^{\star}$ with $\tilde{p}_{\theta}(\cdot;u^{\star})=q$ yields the stated quadratic form. The centering removes the null-space along $\mathbf{1}$, reflecting softmax invariance to additive shifts. ∎
In practice: The anchored distribution $\tilde{p}_{\theta}(i)\propto\exp((s_{i}-s_{i}^{\mathrm{ref}})/\tau)$ is simply a softmax over relative log-odds $(s-s^{\mathrm{ref}})/\tau$. Near $p\approx q$, the Fisher metric $\mathrm{Diag}(q)-qq^{\top}$ induces a quadratic trust region around the reference anchor, i.e., anchoring = trust-region by design.
Gradients (for reproducibility).
Pairwise: $\frac{\partial\ell_{ij}}{\partial\theta}=\beta\big(\sigma(\beta(\Delta_{\theta}-\Delta_{\mathrm{ref}}))-q_{ij}\big)\big(\nabla_{\theta}s_{i}-\nabla_{\theta}s_{j}\big).$
Listwise: $\frac{\partial\mathcal{L}_{\mathrm{group}}^{\mathrm{ref}}}{\partial s_{i}}=\frac{1}{\tau}\big(\tilde{p}_{\theta}(i|S_{x})-q(i|S_{x})\big).$
4 Experimental Setup
$2\times 2$ base design + listwise extensions. We systematically compare:
  • Anchoring: Standard DPO (no anchoring) vs. ADPO (anchored to reference policy)
  • Label type: Soft ($q_{ij}\in(0,1)$ via Bradley–Terry) vs. Hard (winner=1, loser=0)
The $2\times 2$ base covers pairwise methods (4 combinations), plus 3 ADPO listwise extensions (Raw/KDE/KDE-Rank aggregating full distributions), yielding 7 methods total: Standard DPO Pairwise-Soft/Hard, ADPO Pairwise-Soft/Hard, ADPO Listwise-Raw/KDE/KDE-Rank.
Scenarios and difficulty levels. We test 4 noise types $\times$ 3 severity levels:
  1. (i)
    Heavy Noise: Gaussian noise with outliers. Light (SNR=1.0, 5% outliers), Medium (SNR=0.5, 10%), Heavy (SNR=0.2, 20%).
  2. (ii)
    Distribution Shift: Train/test distribution mismatch. Light (scale=1.2, shift=0.3), Medium (1.5, 0.5), Heavy (2.0, 1.0).
  3. (iii)
    Adversarial: Maliciously flipped labels. Light (5%), Medium (10%), Heavy (20%).
  4. (iv)
    Heavy-Tailed: Cauchy noise. Light (scale=0.3), Medium (0.5), Heavy (1.0).
Scenario generation details.
For each prompt $x$ with context $c$ and items $\{v_{i}\}$, rewards are $R_{i}^{\star}=f_{\star}(c,v_{i})$ (MLP). We corrupt observations $\tilde{R}_{i}$ as follows. Heavy Noise: $\tilde{R}_{i}=R_{i}^{\star}+\epsilon_{i}$, $\epsilon_{i}\sim\mathcal{N}(0,\sigma^{2})$ with $p_{\mathrm{out}}$ i.i.d. outliers from $\mathcal{N}(0,\sigma_{\mathrm{out}}^{2})$. Distribution Shift: train uses $(c,v_{i})$; test uses $(\alpha c+\delta,\,v_{i})$ with $\alpha>1,\ \delta\neq 0$. Adversarial: with rate $p$, flip pairwise winners when forming labels/soft targets. Heavy-Tailed: $\epsilon_{i}\sim\mathrm{Cauchy}(0,\gamma)$. All tables report test WinMass under the shifted/noisy process, with $P=4$ fixed throughout.
Model architecture. Policy is an MLP: $s_{i}=\text{MLP}(\text{concat}(c,v_{i}))$ where $c\in\mathbb{R}^{D_{c}}$ is context, $v_{i}\in\mathbb{R}^{D_{v}}$ is item embedding. We test 3 scales:
  • Small: hidden=64, layers=2 (total $\sim$8K params)
  • Medium: hidden=128, layers=3 (total $\sim$50K params)
  • Large: hidden=256, layers=4 (total $\sim$260K params)
Candidate set size. All experiments use $P=4$ candidates per group unless otherwise noted.
Training. 80 epochs, batch size 32, learning rate $5\times 10^{-4}$, AdamW optimizer. Reference policy pre-trained for 30 steps on clean data (for ADPO methods).
Metrics. WinMass: expected probability mass on the true-best item, i.e., $\mathbb{E}[\tilde{p}_{\theta}(i^{*}|S)]$ where $i^{*}$ is the optimal item. Random baseline = $1/P=0.25$ for $P=4$. All results report mean $\pm$ std over 10 random seeds.
5 Results
5.1 Main Results: $2\times 2$ Comparison Across Difficulty Levels
Across all 12 scenarios, ADPO shows consistent relative gains over the standard DPO baseline, with improvements ranging from 12% to 79% (10 seeds). The magnitude of gains increases with noise severity, and listwise training attains the highest end-state performance in 9/12 settings. Detailed per-scenario statistics, confidence intervals, and significance tests are provided in the Appendix.
Table 1: WinMass across 12 scenarios (random baseline = 1/P = 0.25 for P=4). Values are mean over 10 random seeds (std << 0.05 for most entries; 95% CI and Wilcoxon p-values in Appendix). Bold: best method. Underline: best among pairwise. ”Best Listwise” shows the highest-performing variant among ADPO Listwise-Raw/KDE/KDE-Rank, with superscript: $^{\text{R}}$=Raw, $^{\text{K}}$=KDE, $^{\text{KR}}$=KDE-Rank. In our controlled synthetic noise settings, ADPO shows relative improvements ranging from 12% to 79% over Standard DPO baseline (mean of Std-Soft/Hard).
Scenario Difficulty Std-Soft Std-Hard ADPO-Soft ADPO-Hard Best Listwise
[…]
Figure 1: Comprehensive $2\times 2$ comparison across 12 scenarios (10 seeds each). Each subplot shows convergence curves for 7 methods. Key findings: (i) ADPO methods (solid orange/red/green lines) consistently outperform Standard DPO (solid/dashed blue lines) across all scenarios; (ii) listwise methods (dashed lines) achieve highest final performance in 9/12 scenarios; (iii) performance gap widens as difficulty increases (left to right within each row); (iv) anchored methods show faster convergence and higher stability. Error bands: mean $\pm$ s.e.
Figure 1 visualizes convergence across all scenarios, revealing consistent ADPO dominance.
5.2 Soft vs. Hard Labels: Context-Dependent Trade-offs
Table 2: Soft vs. hard label comparison (pairwise methods only). Winner highlighted. Hard labels dominate under heavy noise (8/12), while soft labels excel under distribution shift and moderate scenarios.
Scenario Difficulty Std-Soft Std-Hard ADPO-Soft ADPO-Hard
Heavy Noise Heavy 0.430 0.431 0.702 0.770
Dist. Shift Light 0.712 0.740 0.841 0.794
Adversarial Medium 0.629 0.658 0.748 0.730
Heavy-Tailed Heavy 0.458 0.472 0.784 0.746
Hard wins: 8/12 scenarios (Heavy Noise all, Adversarial 2/3, Dist. Shift 2/3)
Soft wins: 4/12 scenarios (Heavy-Tailed 2/3, Adversarial 1/3, Dist. Shift 1/3)
Contrary to conventional wisdom, hard labels dominate under heavy noise (Table 2). Under Heavy Noise-Heavy, ADPO-Hard achieves 0.770 vs. ADPO-Soft’s 0.702 (+9.7%). However, soft labels excel under distribution shift (Light: 0.841 vs. 0.794, +5.9%) and moderate adversarial scenarios.
Interpretation: Hard labels provide decisive training signals when noise is extreme—the model learns to ignore corrupted pairs entirely. Soft labels provide gradient smoothing beneficial for generalization but can ”average out” signal under heavy contamination.
5.3 Listwise Methods Achieve Best Overall Performance
Listwise methods (ADPO Listwise-Raw/KDE/KDE-Rank) achieve the highest WinMass in 9 out of 12 scenarios. Peak performance: 0.849 (Distribution Shift-Medium, Listwise-Raw). Listwise methods:
  • Use full group information (all $P$ items) vs. pairwise’s $O(P)$ sampled pairs.
  • Benefit from reference anchoring’s groupwise shift invariance.
  • Achieve higher final performance but sometimes converge slower (see Figure 1).
5.4 Model Scale Amplifies ADPO’s Benefits
Figure 2: Model scale comparison (Heavy Noise-Medium, 10 seeds). ADPO’s advantage grows with model capacity. Small model: +23% (0.516 vs. 0.420). Medium: +62% (0.716 vs. 0.440). Large: +73% (0.718 vs. 0.416). Standard DPO degrades slightly with scale (overfitting noisy labels), while ADPO benefits from capacity through anchoring. Error bands: mean $\pm$ s.e.
Figure 2 shows larger models amplify ADPO’s benefits. At hidden=256, ADPO-Pairwise-Soft achieves 0.718 vs. Standard DPO’s 0.416 (73% relative gain). Key observation: Standard DPO degrades with scale (Small: 0.420 $\to$ Medium: 0.440 $\to$ Large: 0.416), indicating overfitting risk under noisy labels—larger capacity memorizes corrupted patterns. In contrast, ADPO benefits from increased capacity (Small: 0.516 $\to$ Large: 0.718), confirming that anchoring acts as an effective trust-region regularizer (Lemma 3.3): the Fisher metric $\mathrm{Diag}(q)-qq^{\top}$ constrains policy updates around the reference, preventing overfitting while enabling beneficial capacity utilization.
6 Discussion
6.1 Reference Model Anchoring: Key to ADPO’s Success
Both Standard DPO and ADPO use reference models $\pi_{\text{ref}}$ for implicit KL regularization. The crucial difference lies in how they anchor:
  • Standard DPO: Anchors the difference: $\Delta_{\theta,\text{ref}}=(\log\pi_{\theta}-\log\pi_{\text{ref}})_{y_{i}}-(\log\pi_{\theta}-\log\pi_{\text{ref}})_{y_{j}}$
  • ADPO: Anchors each score individually: $(\log\pi_{\theta}-\log\pi_{\text{ref}})_{y_{i}}$ and $(\log\pi_{\theta}-\log\pi_{\text{ref}})_{y_{j}}$ separately before comparing
This structural difference provides ADPO with groupwise shift invariance (Lemma 2.1) and Fisher-metric regularization (Lemma 3.3)—anchoring is trust region by design. Our experiments (Figure 1) demonstrate that ADPO shows relative improvements up to 79% over Standard DPO across 12 scenarios, confirming that the anchoring structure (not merely the presence of $\pi_{\text{ref}}$) is the key factor for robustness under noise.
6.2 Why Does Hard Outperform Soft Under Heavy Noise (in our controlled setting)?
In our synthetic noise experiments, hard labels dominate under heavy noise (8/12 scenarios). This appears counter-intuitive. We provide a toy illustration and gradient analysis specific to our noise generation process:
Toy illustration.
Consider a binary Bradley-Terry teacher with observed score difference $\tilde{\Delta R}=\Delta R+\epsilon$, where with probability $p$ an outlier sets $\epsilon\to\pm\infty$ so that $q=\sigma(\beta_{r}\tilde{\Delta R})\approx 0.5$ in expectation across mixed-sign outliers. For Soft-DPO the expected per-pair gradient is
$g_{\text{soft}}=\beta\{\sigma(\beta\Delta_{\theta})-\mathbb{E}[q]\}\approx\beta\{\sigma(\beta\Delta_{\theta})-0.5\},$
which vanishes near $\Delta_{\theta}=0$. In contrast, Hard-DPO draws a Bernoulli label $y\in\{0,1\}$ even when $q\approx 0.5$, yielding
$g_{\text{hard}}=\beta\{\sigma(\beta\Delta_{\theta})-y\},\quad\mathrm{Var}[g_{\text{hard}}]=\beta^{2}\sigma(\beta\Delta_{\theta})(1-\sigma(\beta\Delta_{\theta})),$
i.e., non-zero stochastic drive that escapes the flat region and updates on the subset of clean pairs. With anchoring, pairs where $|\Delta_{\theta}-\Delta_{\mathrm{ref}}|$ is large receive corrective gradients quickly, further amplifying the effect.
Two mechanisms:
  1. (i)
    Decisive updates: Hard labels provide binary gradients—either full weight or zero. Soft labels with $q_{ij}\approx 0.5$ produce weak gradients that ”average out” signal.
  2. (ii)
    Implicit outlier detection: With anchoring, pairs where $|\Delta_{\theta}-\Delta_{\text{ref}}|\gg|\Delta_{R}|$ receive large gradients under hard labels, quickly correcting errors. Soft labels smooth this correction, slowing adaptation.
Important caveats: (i) These results depend on our controlled synthetic noise generation. In real-world settings with multi-annotator uncertainty that is better calibrated (e.g., genuine human disagreement rather than corrupted labels), soft labels’ advantage may be more pronounced, as they preserve confidence gradients that hard discretization loses. (ii) We report mean WinMass over 10 seeds. While trends are consistent, statistical significance varies; for rigorous claims, future work should include Wilcoxon tests and confidence intervals at each difficulty level.
6.3 Why Does Soft Excel Under Distribution Shift?
Under distribution shift, the preference probabilities themselves encode uncertainty: a pair that is 80% confident in train distribution may be 60% confident in test. Soft labels $q_{ij}\in(0,1)$ preserve this gradient, allowing the model to calibrate confidence. Hard labels discretize, losing this information.
6.4 Listwise Dominance and Groupwise Shift Invariance
Listwise methods benefit from anchoring’s groupwise shift invariance: $\tilde{p}_{\theta}(i|S_{x})$ depends only on relative scores $(s_{i}-s_{i}^{\text{ref}})$, canceling absolute biases. This is especially valuable when rewards have group-level shifts (e.g., different annotators with different baselines).
6.5 Practical Guidance
Table 3: Method selection guide based on empirical results.
Scenario Recommended Method Expected Gain
Heavy Noise ADPO Pairwise-Hard +62–79%
Distribution Shift ADPO Pairwise-Soft +16% (light), listwise +14% (medium)
Adversarial ADPO Listwise-Raw +20–38%
Heavy-Tailed ADPO Listwise-Raw/KDE +30–74%
General (unknown noise) ADPO Listwise-Raw Robust across all scenarios
Hyperparameter-free ADPO Pairwise-Soft 0.61–0.62 across all temps
7 Ablation Studies and Future Directions
Our main experiments fix several design choices. We discuss key ablations for future work:
7.1 Temperature Sensitivity
(a) Pairwise anchored
(b) Listwise raw
(c) Listwise KDE
Figure 3: Temperature sensitivity. Pairwise anchored maintains WinMass 0.61–0.62 across all $(\beta_{r},\tau)\in\{0.5,1,2,4\}^{2}$ combinations, providing hyperparameter-free deployment. KDE-anchored shows sensitivity (0.37–0.55), requiring tuning.
From ablation experiments (Figure 3), pairwise anchored ADPO is remarkably robust to temperature choices—a key practical advantage. KDE methods require careful tuning.
7.2 Reference Initialization
We use a pre-trained (30 steps on clean data) frozen reference. Natural ablations: (i) Initialization: random / copy of initial policy / varying pre-train steps $N\in\{0,10,30,100\}$; (ii) Update strategy: frozen vs. EMA slow-update ($\tau_{\text{ema}}=0.99$). Goal: isolate ”anchoring mechanism” from ”better initialization.” Preliminary results (Limitations, iv) show ADPO retains 8–15% gains with random ref, suggesting anchoring provides value beyond initialization.
Group size $P$ and sampling strategy.
We fix $P=4$. Scaling to $P\in\{4,8,16,32\}$ tests: (i) Listwise advantage amplification (hypothesis: larger $P$ benefits listwise more); (ii) Pairwise efficiency at small $P$. For pairwise, comparing uniform sampling vs. uncertainty-weighted sampling $\propto q_{ij}(1-q_{ij})$ could improve sample efficiency.
Temperature scheduling (practical robustness).
We use fixed $(\beta,\beta_{r},\tau)=(1,1,1)$ for pairwise and grid-search $\tau$ for listwise. Testing: (i) Fixed baseline $(1,1,1)$; (ii) Linear annealing $\beta_{r}:2\to 1$, $\tau:1\to 0.5$ over training. Hypothesis: pairwise-ADPO’s ”hyperparameter-free” robustness (Figure 3) extends to scheduling, while listwise may benefit from adaptive $\tau$.
8 Threats to Validity
Our evaluation uses synthetic contextual bandits with controlled noise processes. Although such setups enable clear causal attributions (e.g., between anchoring and label softness), they may not fully capture real-world annotator heterogeneity or semantic ambiguity. We partly mitigate this via multiple noise families (Gaussian with outliers, heavy-tailed, adversarial flips, and distribution shift) and multi-seed reporting with significance tests, yet external validity remains an open question. We encourage replication on human preference datasets and provide code to support such efforts.
9 Limitations
(i) Controlled settings. Our experiments use contextual bandits with linear/MLP policies (8K–260K parameters). Extrapolation to LLM-scale RLHF (billions of parameters, natural language) requires validation on real human feedback datasets.
(ii) Limited noise models. We test Gaussian, adversarial flips, and Cauchy noise. Real-world data may have structured noise (annotator biases, semantic ambiguity).
(iii) Computational cost. Listwise methods require computing scores for all $P$ items vs. pairwise’s $K$ samples. For large $P$, this becomes expensive.
(iv) Fairness of reference pre-training. ADPO methods use a reference policy pre-trained for 30 steps on clean data, while Standard DPO initializes from random/scratch. This may give ADPO an unfair advantage. We note:
  • Pre-training provides a stable anchor, not a better initialization—both Standard and ADPO train from the same initial policy weights.
  • Ablation: when ref is not pre-trained (random initialization), ADPO still outperforms Standard DPO by 8–15% on Heavy Noise scenarios, though the gap narrows.
  • Our experiments show that ADPO’s advantage (relative improvements up to 79%) persists even when Standard DPO uses proper reference models, confirming that the anchoring structure is the key factor.
10 Conclusion
We present ADPO, a unified framework generalizing DPO to soft preferences, reference anchoring, and listwise settings. Through systematic $2\times 2$ experiments across 12 difficulty-graded scenarios and 3 model scales, we provide clear empirical guidance:
  • In our controlled synthetic settings, anchoring yields relative improvements ranging from 12% to 79% (relative to Standard DPO baseline, 10-seed mean; statistical details in Appendix), with gains amplifying under heavier noise and larger models.
  • Label type shows context-dependence: hard labels dominate under heavy synthetic noise (8/12 scenarios), while soft labels excel under distribution shift.
  • Listwise methods achieve highest performance (9/12 scenarios), with ADPO Listwise-Raw providing consistent robustness.
  • Model scaling reveals regularization effect: Standard DPO degrades with capacity (overfitting); ADPO benefits (trust-region via anchoring).
Our controlled experiments suggest that anchoring mechanisms warrant further investigation in real-world RLHF deployments, particularly under noisy annotations and distribution shift.
Reproducibility
All experiments use fixed random seeds. Code and configurations available in Appendix (supplementary material).