### v6

# ADPO: Anchored Direct Preference Optimization A Unified Framework from Pairwise to Listwise Preferences

Abstract
[…]
1 Introduction
Preference optimization has become the dominant paradigm for aligning large language models with human values, with Direct Preference Optimization (DPO) [1] emerging as a simple and effective alternative to traditional reinforcement learning from human feedback (RLHF) [2, 3]. However, DPO’s reliance on hard, pairwise preference labels makes it brittle. In noisy settings, this fragility can be severe; our controlled experiments (Section 5) show that under heavy noise, DPO’s performance degrades by up to 93% relative to our method (0.375 vs. 0.725 WinMass). This sensitivity arises because the DPO objective only regularizes differences of log-probabilities, providing a weak inductive bias that is susceptible to overfitting corrupted labels.
We introduce Anchored Direct Preference Optimization (ADPO), a unified framework that connects seemingly disparate learning paradigms. Through an anchored projection—minimizing $\mathrm{KL}\!\big(q\,\|\,\tilde{p}_{\theta}\big)$ where $\tilde{p}_{\theta}$ is defined by anchored logits $(\ell-\ell^{\mathrm{ref}})/\tau_{\mathrm{anc}}$—we show that supervised fine-tuning, knowledge distillation, maximum-entropy RL policy updates, advantage-weighted methods (AWR/AWAC/MPO), and DPO all arise from different choices of the target distribution $q$, the anchor policy $\pi_{\mathrm{ref}}$, and the temperature $\tau_{\mathrm{anc}}$ (Section 3.2, Proposition 3.4). By centering the coordinate system at a reference policy before applying softmax, ADPO provides groupwise shift invariance and Fisher-metric regularization—properties absent in standard methods.
We further show that this anchoring induces a principled geometric structure: the softmax Fisher information metric governs the local curvature around the target distribution (Lemma 3.6), yielding an implicit trust region centered at the target (not at the anchor). This offers a distribution-space analogue to TRPO [23] without explicit constraints. Because the geometry depends only on $q$, we can cleanly compare anchor update strategies. Under our experimental settings we find a task-dependent trade-off:
  • Dynamic anchors suit online exploration in noisy environments, improving performance by 5–11% in 3/4 scenarios on our benchmarks.
[…]
2 Related Work
Preference optimization. More generally, the Preference Optimization (PO) framework [18] formulates an entropy-regularized objective aligning policy probabilities with qualitative preference signals, which inspires our entropy-regularized interpretation of anchoring (Section 3.2), though our work focuses on preference optimization for RLHF-style alignment rather than combinatorial optimization. DPO [1] reparameterizes the reward model through policy-reference log-ratios, enabling direct optimization. Extensions include identity-free preferences (IPO [5]), contrastive objectives (CPO [6]), and iterative refinement [7]. Our work provides a unified framework encompassing these variants.
Listwise preference learning. Plackett–Luce models [9, 10] enable listwise preferences through recursive top-1 selections. Recent work applies PL to LLM alignment [11, 12, 13]. We extend DPO to listwise settings through reference anchoring.
Robust learning under noise. Prior work addresses noise through majority voting [15], uncertainty quantification [16], and robust reward learning [17]. Our approach encodes uncertainty in soft probabilities and provides anchoring for groupwise shift invariance.
Relation to KD and TRPO.
Conventional knowledge distillation (KD) [22] aligns a student with a teacher by minimizing $\mathrm{KL}(q\|p_{\theta})$, which directly matches absolute probabilities in the (unanchored) probability simplex. While effective, this objective lacks a fixed geometric center and thus provides no notion of a trust region, making it sensitive to teacher noise and distributional shifts. In contrast, trust-region policy optimization (TRPO) [23] explicitly constrains $\mathrm{KL}(\pi_{\theta_{\text{old}}}\|\pi_{\theta})$ in the parameter space, defining a moving local quadratic region that stabilizes updates.
Our ADPO formulation can be viewed as a distribution-space analogue of TRPO: by anchoring logits to a reference policy, the optimization is performed in coordinates centered at the reference, while the local quadratic well (Fisher metric) is centered at the target distribution $q$, yielding an implicit trust region without an explicit KL constraint. To our knowledge, a direct anchored cross-entropy formulation emphasizing a fixed geometric center in probability space has not been formalized in prior work, though related ideas appear in advantage-weighted methods [24, 25] and maximum a posteriori policy optimization [26].
Entropy-regularized grounding and its link to ADPO.
Under the maximum-entropy RL objective
$\max_{\pi}\ \mathbb{E}_{s\sim\mathcal{D}}\Big[\mathbb{E}_{a\sim\pi(\cdot|s)}r(s,a)\ +\ \alpha\,\mathsf{H}\big(\pi(\cdot|s)\big)\Big],$
the optimal policy admits the Boltzmann form [19] (also known as energy-based policies [21])
$\pi^{*}(a|s)\;=\;\frac{1}{Z(s)}\exp\!\Big(\tfrac{1}{\alpha}\,r(s,a)\Big),$ (1)
and the reward can be reparameterized by the (optimal) policy log-probability
$r(s,a)\;=\;\alpha\,\log\pi^{*}(a|s)\;+\;\alpha\,\log Z(s),$ (2)
so any relative quantity (pairwise/listwise) cancels $Z(s)$. Eqs. (1)–(2) yield a direct bridge between reward differences and policy log-odds: Bradley–Terry and Plackett–Luce targets arise by mapping $\Delta r$ to a preference probability via a sigmoid/softmax, while ADPO matches the student’s anchored log-odds $(\ell-\ell^{\mathrm{ref}})/\tau_{\mathrm{anc}}$ to these soft targets. A full derivation of (1) as the minimizer of $\mathbb{E}_{s}\,\mathrm{KL}\big(\pi(\cdot|s)\,\|\,\pi^{*}(\cdot|s)\big)$ and the log-partition cancellation in pairwise/listwise forms can be found in [18] (see their Eqs. (3)–(6), (20)–(27)).
Lemma 2.1 (Groupwise shift invariance).
For any context-independent shift $h(s)$, if $r^{\prime}(s,a)=r(s,a)-h(s)$ then $\pi^{*}$ in (1) is unchanged; hence all pairwise/listwise probabilities (BT/PL) are identical, and anchored ADPO—which operates on $(\ell-\ell^{\mathrm{ref}})$—inherits groupwise shift invariance.
Proof sketch.
For any group-constant $c$, $\mathrm{softmax}(u)=\mathrm{softmax}(u+c\mathbf{1})$ by shift-invariance. Since $r^{\prime}(s,a)=r(s,a)-h(s)$ only changes the group constant (log-partition $\log Z(s)$ in Eq. (2)), the Bradley–Terry/Plackett–Luce probabilities and anchored distribution $\tilde{p}$ remain unchanged. See [18], Prop. 3.1 and App. D.2 for full derivation. ∎
3 Anchored Direct Preference Optimization (ADPO)
3.1 Problem Setting and Anchored Softmax
Motivation: Why anchoring?
Traditional preference optimization methods (e.g., DPO with uniform reference) operate on absolute log-probabilities $\ell_{i}=\log\pi_{\theta}(y_{i}|x)$, making them sensitive to:
  • Annotator baseline biases: Different annotators may have systematically different baseline preferences
  • Sequence-level biases: Length bias in language models, state-value offsets in RL
  • Distribution shift: Changes in the reference distribution between training and deployment
ADPO addresses these issues by introducing a reference anchor $\pi_{\mathrm{ref}}$ and working with relative log-probabilities $(\ell-\ell^{\mathrm{ref}})$, which automatically cancels group-level baseline shifts through the softmax’s shift-invariance property (detailed below).
Notation.
Let $x\!\sim\!\mathcal{D}$ be a context (e.g., prompt or state). For a candidate set $S_{x}=\{y_{1},\dots,y_{P}\}$ (e.g., responses or actions), we define:
  • Student policy $\pi_{\theta}(y|x)$ with log-probabilities $\ell_{i}=\log\pi_{\theta}(y_{i}|x)$
  • Anchor policy $\pi_{\mathrm{ref}}(y|x)$ with log-probabilities $\ell_{i}^{\mathrm{ref}}=\log\pi_{\mathrm{ref}}(y_{i}|x)$
  • Target preference $q(\cdot|S_{x})$ with $q\succ 0$ (strictly positive distribution)
We work with anchored logits:
$u_{i}\;=\;\frac{\ell_{i}-\ell_{i}^{\mathrm{ref}}}{\tau_{\mathrm{anc}}}\,,\qquad\bar{u}_{i}\;=\;u_{i}-\sum_{j\in S_{x}}q(j|S_{x})\,u_{j},$ (3)
where $\tau_{\mathrm{anc}}>0$ is an anchoring temperature controlling regularization strength, and $\bar{u}$ denotes the $q$-centered anchored logits (always applied for numerical stability and geometric consistency).
Groupwise shift invariance.
Because $\tilde{p}_{\theta}$ depends only on differences of anchored logits $u_{i}-u_{j}$, adding a constant to all $(\ell_{i}-\ell_{i}^{\mathrm{ref}})$ within a group leaves $\tilde{p}_{\theta}$ unchanged. This property automatically cancels group-level baselines such as:
  • Annotator-specific biases (different annotators may have different baseline preferences)
  • Sequence-level biases (e.g., length bias in language models)
  • State-dependent baselines in RL (e.g., value function offsets)
Remark 3.1 (Sequence tasks and length normalization).
For sequence generation tasks, one typically constructs $\ell$ and $\ell^{\mathrm{ref}}$ using length-normalized log-probabilities to avoid length bias. The affine invariance of the groupwise softmax ensures all results remain valid.
Target distribution $q$.
The target $q(\cdot|S_{x})$ encodes the desired preference over candidates within group $S_{x}$. When the full action/candidate space is considered, we write $q(\cdot|x)$; the in-group distribution can be viewed as conditioning on $S_{x}$. Common choices include:
  • Plackett–Luce (PL) soft labels from rankings
  • Bradley–Terry (BT) soft labels from pairwise comparisons: $q(i\succ j|S_{x})=\sigma(\beta_{r}\Delta R_{ij})$
  • Q-function based: $q(a|s)\propto\exp(Q(s,a)/\beta_{r})$ in RL settings
  • Teacher distribution: $q=\pi_{\text{teacher}}$ for knowledge distillation
3.2 Unified ADPO Objective and Its Special Cases
Our master objective is the groupwise forward KL divergence:
$\boxed{\min_{\theta}\ \mathcal{L}_{\mathrm{ADPO}}\;=\;\mathbb{E}_{x,S_{x}}\!\Big[\mathrm{KL}\!\big(q(\cdot|S_{x})\,\big\|\,\tilde{p}_{\theta}(\cdot|S_{x})\big)\Big],\quad\tilde{p}_{\theta}=\mathrm{softmax}(\bar{u}),\ \bar{u}=\tfrac{(\ell-\ell^{\mathrm{ref}})}{\tau_{\mathrm{anc}}}-\langle q,\tfrac{(\ell-\ell^{\mathrm{ref}})}{\tau_{\mathrm{anc}}}\rangle\mathbf{1}.}$ (4)
(Equivalently, using $u$ without centering yields the same $\tilde{p}$ up to a constant shift, by softmax shift-invariance.)
Why forward KL?
Forward KL $\mathrm{KL}(q\|\tilde{p}_{\theta})$ is mode-covering (encourages $\tilde{p}_{\theta}$ to place mass wherever $q$ does), making it robust to soft/noisy targets; reverse KL is mode-seeking (concentrates on a single mode), which is brittle under label noise. For listwise supervision with uncertain preferences, mode-covering is more stable.
Within a group, this equals a cross-entropy up to a constant. Writing $\mathcal{L}(u)=A(u)-\langle q,u\rangle$ where $A(u)=\log\sum_{j}e^{u_{j}}$ is the log-partition function, the loss is strictly convex in $u$ (on $\mathbf{1}^{\perp}$):
$\sum_{i}q_{i}\log\tilde{p}_{\theta,i}\;=\;-\frac{1}{\tau_{\mathrm{anc}}}\sum_{i}q_{i}\big(\ell_{i}-\ell_{i}^{\mathrm{ref}}\big)\;+\;\log\!\sum_{j}\exp\!\Big(\frac{\ell_{j}-\ell_{j}^{\mathrm{ref}}}{\tau_{\mathrm{anc}}}\Big)\;=\;\mathrm{KL}(q\|\tilde{p}_{\theta})+H(q).$ (5)
Closed-form optimum in distribution space.
If the model class is expressive enough, the optimum of (4) satisfies:
Proposition 3.2 (Closed-form optimum).
When $\tilde{p}_{\theta}(\cdot|S_{x})=q(\cdot|S_{x})$ is attainable, the optimal policy satisfies
$\pi_{\theta}(y|x)\;\propto\;\pi_{\mathrm{ref}}(y|x)\,q(y|x)^{\tau_{\mathrm{anc}}}.$ (6)
This requires mild support conditions: in the discrete case, $q(y|x)>0\Rightarrow\pi_{\mathrm{ref}}(y|x)>0$; in the continuous case, absolute continuity and integrability ensure $\tilde{p}=q$ is attainable and the proportionality is normalizable (well-defined). When training on sampled groups $S_{x}$, this gives the conditional optimum within each group; if groups are uniformly sampled from the full action space and the model is sufficiently expressive, the empirical optimum converges to the population optimum.
Proof sketch.
At optimum, $\tilde{p}_{\theta}(i|S_{x})=q(i|S_{x})$. By definition,
$\tilde{p}_{\theta}(i|S_{x})=\mathrm{softmax}\!\Big(\frac{\ell_{i}-\ell_{i}^{\mathrm{ref}}}{\tau_{\mathrm{anc}}}\Big)\propto\exp\!\Big(\frac{\ell_{i}-\ell_{i}^{\mathrm{ref}}}{\tau_{\mathrm{anc}}}\Big)=\Big(\frac{\pi_{\theta}(i|x)}{\pi_{\mathrm{ref}}(i|x)}\Big)^{1/\tau_{\mathrm{anc}}}.$
Setting this equal to $q(i|S_{x})$ and solving yields (6). ∎
Temperature interpretation.
  • Small $\tau_{\mathrm{anc}}$: Strong regularization toward anchor $\pi_{\mathrm{ref}}$ (conservative updates). The local curvature scales as $1/\tau_{\mathrm{anc}}^{2}$ (Section 3.3), yielding a tighter implicit trust region.
  • Large $\tau_{\mathrm{anc}}$: Weak regularization, $\pi_{\theta}$ tracks $q$ more closely (aggressive updates)
  • $\tau_{\mathrm{anc}}=1$: Balanced regime (commonly used in practice)
Pairwise and DPO as corollaries.
For $|S_{x}|{=}2$, let $\Delta_{\theta}=\ell_{i}-\ell_{j}$ and $\Delta_{\mathrm{ref}}=\ell_{i}^{\mathrm{ref}}-\ell_{j}^{\mathrm{ref}}$. With a Bradley–Terry target $q_{ij}\in(0,1)$, (4) reduces to:
$\boxed{\mathcal{L}_{ij}^{\text{ADPO}}=\log\!\big(1+\exp[(\Delta_{\theta}-\Delta_{\mathrm{ref}})/\tau_{\mathrm{anc}}]\big)-q_{ij}\,(\Delta_{\theta}-\Delta_{\mathrm{ref}})/\tau_{\mathrm{anc}},}$ (7)
which recovers the standard DPO loss [1] when $\Delta_{\mathrm{ref}}=0$ (uniform anchor) and $q_{ij}\in\{0,1\}$ (hard labels). When $q_{ij}\in(0,1)$ (soft labels from noisy annotations), this becomes "soft DPO" that naturally handles label uncertainty. When $\Delta_{\mathrm{ref}}\neq 0$ (non-uniform anchor), we obtain "anchored DPO" with improved stability.
Gradient.
The gradient of the pairwise loss is:
$\frac{\partial\mathcal{L}_{ij}^{\mathrm{ADPO}}}{\partial\theta}=\frac{1}{\tau_{\mathrm{anc}}}\big(\sigma[(\Delta_{\theta}-\Delta_{\mathrm{ref}})/\tau_{\mathrm{anc}}]-q_{ij}\big)(\nabla_{\theta}\ell_{i}-\nabla_{\theta}\ell_{j}),$
where $\sigma(\cdot)$ is the sigmoid function.
Listwise (general) case.
For candidate sets $S_{x}=\{y_{1},\dots,y_{P}\}$ with $P>2$, the ADPO listwise objective is:
$\boxed{\mathcal{L}_{\text{group}}^{\text{ADPO}}=-\sum_{i\in S_{x}}q(i|S_{x})\log\tilde{p}_{\theta}(i|S_{x}),\qquad\tilde{p}_{\theta}(\cdot|S_{x})=\mathrm{softmax}(\bar{u}),}$ (8)
with gradient
$\frac{\partial\mathcal{L}_{\mathrm{group}}^{\text{ADPO}}}{\partial\ell_{i}}=\frac{1}{\tau_{\mathrm{anc}}}\big(\tilde{p}_{\theta}(i|S_{x})-q(i|S_{x})\big).$ (9)
Setting P=2P{=}2 recovers the binary case (7), showing that ADPO continuously extends DPO from pairwise to listwise soft supervision.
$q$-centering: A principled use of shift-invariance.
Recall from Section 3.1 that softmax is shift-invariant: $\mathrm{softmax}(u)=\mathrm{softmax}(u+c\mathbf{1})$ for any constant $c$. We exploit this property by choosing $c=-\sum_{j}q_{j}u_{j}$ (the negative $q$-weighted mean), yielding:
$\boxed{\bar{u}_{i}=u_{i}-\sum_{j\in S_{x}}q(j|S_{x})\,u_{j},\qquad u_{i}=\frac{\ell_{i}-\ell_{i}^{\mathrm{ref}}}{\tau_{\mathrm{anc}}}.}$ (10)
This $q$-centering is always applied in our algorithm (Algorithm 1, both modes) and serves three purposes: numerical stability (centers logits near zero); geometric consistency (removes null-space along $\mathbf{1}$ in Fisher metric, Section 3.3); theoretical alignment (consistent with $q$-centered deviation $\bar{\delta}$ in Lemma 3.6).
Our final anchored distribution is thus:
$\tilde{p}_{\theta}(i|S_{x})=\mathrm{softmax}(\bar{u})_{i}=\frac{\exp(\bar{u}_{i})}{\sum_{j\in S_{x}}\exp(\bar{u}_{j})}.$ (11)
Remark 3.3 (Optional: Fisher variance scaling).
For extremely large $P$ (e.g., $P\geq 16$), one can additionally apply Fisher normalization:
$\widehat{u}_{i}=\frac{\bar{u}_{i}}{\sqrt{\mathrm{Var}_{q}[u]+\varepsilon}},\qquad\mathrm{Var}_{q}[u]=\sum_{j}q_{j}u_{j}^{2}-\Big(\sum_{j}q_{j}u_{j}\Big)^{2}.$ (12)
This "whitens" the gradient but changes the effective temperature. We do not use this in our experiments and recommend it only as a troubleshooting step for very large $P$.
Unifying major learning paradigms.
Different choices of $(q,\pi_{\mathrm{ref}},\tau_{\mathrm{anc}})$ recover major learning paradigms:
Proposition 3.4 (ADPO unifies common paradigms).
Specific choices of $(q,\pi_{\mathrm{ref}},\tau_{\mathrm{anc}})$ recover:
  • Supervised Fine-Tuning (SFT): $q$ from labeled data, $\pi_{\mathrm{ref}}$ uniform, $\tau_{\mathrm{anc}}=1$.
  • Knowledge Distillation (KD): $q=\mathrm{softmax}(z_{T}/T)$ from teacher logits $z_{T}$, $\pi_{\mathrm{ref}}$ uniform, $\tau_{\mathrm{anc}}=T$.
  • Maximum-Entropy RL: $q(a|s)\propto\exp(Q/\alpha)$ from Q-values, $\pi_{\mathrm{ref}}$ uniform, $\tau_{\mathrm{anc}}=1$.
  • Direct Preference Optimization (DPO): $|S_{x}|=2$, $q(i)=\sigma(\beta_{r}\Delta R)$ from reward differences, $\pi_{\mathrm{ref}}$ uniform.
Verification.
Each case follows from Proposition 3.2 by substituting the specified $(q,\pi_{\mathrm{ref}},\tau_{\mathrm{anc}})$. ∎
Table 1 summarizes this unification.
Table 1: ADPO as a unifying projection framework. Different choices of target $q$, anchor $\pi_{\mathrm{ref}}$, and temperature $\tau_{\mathrm{anc}}$ recover major learning paradigms. Key: Optimization matches the anchored distribution $\tilde{p}_{\theta}=\mathrm{softmax}((\ell-\ell^{\mathrm{ref}})/\tau_{\mathrm{anc}})$ to the target $q$; the corresponding policy $\pi_{\theta}$ satisfies $\pi_{\theta}\propto\pi_{\mathrm{ref}}\cdot q^{\tau_{\mathrm{anc}}}$ when the optimum is reachable (Prop. 3.2). Caution: When anchor and target both equal the teacher ($\pi_{\mathrm{ref}}=\pi_{T},q=\pi_{T}$), the optimum becomes $\pi_{\theta}\propto\pi_{T}^{1+\tau_{\mathrm{anc}}}$, which sharpens the teacher distribution and may degrade calibration; use small $\tau_{\mathrm{anc}}$ or add auxiliary regularization if needed.
Task Target $q(\cdot\mid x)$ Anchor $\pi_{\mathrm{ref}}$ $\tau_{\mathrm{anc}}$ Recovers
SFT data (one-hot/soft) uniform 11 cross-entropy
KD (standard) $\mathrm{softmax}(z_{T}/T)$ uniform $T$ $T^{2}\,\mathrm{CE}(q_{T},\mathrm{softmax}(\ell/T))$
KD (anchored) teacher $\pi_{T}$ $\pi_{T}$ any teacher-anchored KD
Max-ent RL $\propto\exp(Q/\alpha)$ uniform 11 $\pi_{\theta}\propto\exp(Q/\alpha)$
DPO (binary) Bradley–Terry / hard uniform any original DPO
ADPO (listwise) Plackett–Luce / soft any any this work
Warning: Sharpens teacher; see caption.
3.3 Fisher Geometry: An Implicit Trust Region
The anchored softmax structure induces a natural Riemannian geometry on the probability simplex. Let $\tilde{p}_{\theta}$ (hereafter $p$ for brevity) denote the current anchored distribution. The softmax Fisher information matrix is:
$F\;=\;\operatorname{Diag}(p)-pp^{\top}\;\in\;\mathbb{R}^{P\times P},\qquad F\mathbf{1}{=}0,$ (13)
where $F\mathbf{1}=0$ reflects the normalization constraint and shift-invariance. Near the optimum $u^{\star}$ (where $\tilde{p}_{\theta}(u^{\star})=q$), the KL divergence admits a second-order expansion:
$\mathrm{KL}(q\|\tilde{p}(u))\;\approx\;\tfrac{1}{2\tau_{\mathrm{anc}}^{2}}\,\bar{\delta}^{\top}(\operatorname{Diag}(q)-qq^{\top})\bar{\delta},\qquad\bar{\delta}=(u-u^{\star})-\langle q,u-u^{\star}\rangle\mathbf{1},$ (14)
where the $1/\tau_{\mathrm{anc}}^{2}$ scaling governs the local curvature (see Figure 1). This creates an implicit trust region centered at the target $q$, independent of the anchor $\pi_{\mathrm{ref}}$—the anchor defines the coordinate system, while the Fisher metric determines the geometry.
Figure 1: Fisher geometry as implicit trust region. The Fisher metric $\operatorname{Diag}(q)-qq^{\top}$ defines a quadratic bowl in anchored coordinates $u=(\ell-\ell^{\mathrm{ref}})/\tau_{\mathrm{anc}}$, centered at target $q$. Smaller $\tau_{\mathrm{anc}}$ yields tighter curvature ($\propto 1/\tau_{\mathrm{anc}}^{2}$).
Natural gradient and pullback metric.
The natural gradient $\delta u_{\text{nat}}=F^{+}(q-p)$ (where $F^{+}$ is the Moore–Penrose pseudoinverse) gives the steepest descent direction in the Fisher metric [27]. In parameter space, the pullback metric is:
$\mathsf{G}(\theta;S_{x})=J\,F\,J^{\top}=\sum_{i}p_{i}g_{i}g_{i}^{\top}-\Big(\sum_{i}p_{i}g_{i}\Big)\Big(\sum_{i}p_{i}g_{i}\Big)^{\top},\quad J=[g_{1},\ldots,g_{P}],\ g_{i}=\nabla_{\theta}\ell_{i}.$ (15)
This is exactly the Fisher information matrix used in TRPO [23], but ADPO induces it implicitly through the anchored softmax—no explicit second-order optimization is required.
Remark 3.5 (Theory vs practice).
Our implementation (Algorithm 1, both modes) uses standard first-order optimization (Adam) on the ADPO loss (4). The Fisher geometry provides theoretical insight into why this works well, not how we implement it.
3.4 Mirror-Descent Equivalence and Convergence
The ADPO update can be viewed as mirror descent with KL Bregman divergence. Within each group, the loss $\mathcal{L}(u)=A(u)-\langle q,u\rangle$ (where $A(u)=\log\sum_{j}e^{u_{j}}$) is convex in anchored logits $u=(\ell-\ell^{\mathrm{ref}})/\tau_{\mathrm{anc}}$.
Lemma 3.6 (Implicit trust region).
Near the optimum $u^{\star}$ (where $\tilde{p}_{\theta}(u^{\star})=q$), writing $\bar{\delta}=(u-u^{\star})-\langle q,u-u^{\star}\rangle\mathbf{1}$,
$\mathrm{KL}\big(q\|\tilde{p}(u)\big)\;=\;\tfrac{1}{2\tau_{\mathrm{anc}}^{2}}\,\bar{\delta}^{\top}\big(\operatorname{Diag}(q)-qq^{\top}\big)\bar{\delta}\;+\;o(\|\bar{\delta}\|^{2}).$ (16)
Proof sketch.
Second-order Taylor expansion of $\mathcal{L}$ at $u^{\star}$ with $\nabla^{2}A(u)=\operatorname{Diag}(\tilde{p})-\tilde{p}\,\tilde{p}^{\top}$ yields the stated form. The $q$-centering removes the null-space along $\mathbf{1}$. ∎
Theorem 3.7 (Mirror-descent equivalence).
The ADPO update with step size $\eta$ satisfies:
$u_{t+1}=\arg\min_{u}\big\langle\nabla\mathcal{L}(u_{t}),\,u\big\rangle+\frac{1}{\eta}\,\mathrm{KL}\!\big(\tilde{p}(u_{t})\,\|\,\tilde{p}(u)\big),$
inducing the multiplicative rule $\tilde{p}_{t+1}\propto\tilde{p}_{t}\odot\exp(-\eta(\tilde{p}_{t}-q))$ on distributions. See Appendix for full derivation.
Corollary 3.8 (Dynamic anchor as proximal trust region).
When $\pi_{\mathrm{ref}}=\pi_{\text{old}}$ (dynamic anchor), each step controls $\mathrm{KL}(\tilde{p}_{\text{old}}\|\tilde{p}_{\text{new}})$ in anchored coordinates, mirroring TRPO’s $\mathrm{KL}(\pi_{\text{old}}\|\pi_{\text{new}})$ constraint with groupwise shift invariance. The temperature $\tau_{\mathrm{anc}}$ scales curvature as $1/\tau_{\mathrm{anc}}^{2}$, acting as a radius knob.
3.5 Anchor Strategies: Offline vs Online
The flexibility of ADPO’s anchoring mechanism enables different instantiations depending on the choice of anchor policy $\pi_{\mathrm{ref}}$. Because the Fisher metric depends only on the target $q$ (not on $\pi_{\mathrm{ref}}$), we can cleanly separate the choice of anchor from the learning objective, leading to two main paradigms.
Offline ADPO: Fixed or no anchor.
In offline settings (SFT, KD, preference learning from fixed datasets), the anchor is either:
  • No anchor (uniform): Set $\pi_{\mathrm{ref}}$ uniform, equivalent to $\ell^{\mathrm{ref}}_{i}=0$. This recovers:
    • SFT: $q$ from labeled data, $\tau_{\mathrm{anc}}=1$ → standard cross-entropy
    • KD: $q=\mathrm{softmax}(z_{T}/T)$ from teacher logits, $\tau_{\mathrm{anc}}=T$ → temperature-scaled KD
  • Fixed anchor: Set $\pi_{\mathrm{ref}}$ once (e.g., pretrained model, SFT checkpoint) and never update. The coordinate system remains fixed, providing a stable geometric center for:
    • Aligning to labels/rewards/preferences
    • Regularizing toward a known good policy
    • Preventing catastrophic forgetting
Key property: With fixed/no anchor, the optimization has monotonic convergence guarantees (Theorem 3.10).
Online ADPO: Dynamic anchor.
In online settings (RL, iterative RLHF, continual learning), the anchor is dynamically updated to track the evolving policy:
$\boxed{\mathcal{L}_{\text{ADPO-online}}\;=\;\mathbb{E}_{x}\Big[\mathrm{KL}\!\Big(q(\cdot|x)\ \Big\|\ \mathrm{softmax}\!\big((\ell_{\theta}-\ell_{\text{old}})/\tau_{\mathrm{anc}}\big)\Big)\Big],\quad\pi_{\mathrm{ref}}\leftarrow\pi_{\text{old}}.}$ (17)
(The anchored logits $(\ell_{\theta}-\ell_{\text{old}})/\tau_{\mathrm{anc}}$ are $q$-centered as in Eq. 10.)
Anchor update strategies:
  • Per-iteration: $\pi_{\mathrm{ref}}\leftarrow\pi_{\theta}$ every iteration (like TRPO [23]/PPO [4])
  • Periodic: $\pi_{\mathrm{ref}}\leftarrow\pi_{\theta}$ every $N$ iterations (balances stability/adaptivity)
  • EMA: $\pi_{\mathrm{ref}}\leftarrow\alpha\pi_{\mathrm{ref}}+(1-\alpha)\pi_{\theta}$ (smooth evolution)
Key property: The anchoring defines a moving coordinate system, while the Fisher metric (determined by $q$) provides an implicit trust region. This mirrors TRPO’s $\mathrm{KL}(\pi_{\text{old}}\|\pi_{\text{new}})$ constraint [23] but operates in distribution space with groupwise shift invariance.
Task-dependent trade-off.
Our experiments (Section 5) reveal:
  • Dynamic anchors improve online exploration in noisy environments by 5–11% in 3/4 scenarios (Section 5.4). The adaptive coordinate frame enables effective exploration while maintaining stability.
[…]
Remark 3.9 (Connection to existing methods).
Offline: No anchor recovers SFT/KD; fixed anchor adds regularization toward a reference policy. Online: Dynamic anchor connects to TRPO/PPO but with implicit Fisher-metric trust regions and groupwise shift invariance.
3.6 Stability and Convergence Analysis
Having established the geometric structure (Section 3.3) and algorithmic view (Section 3.4) of ADPO, we now analyze its stability and convergence properties.
Fixed anchor: Strong convergence guarantees.
When the anchor $\pi_{\mathrm{ref}}$ is fixed throughout training, ADPO enjoys strong convexity and linear convergence within the anchored coordinate system.
Theorem 3.10 (Monotonic descent under fixed anchor).
With a fixed anchor, $\mathcal{L}(u)=A(u)-\langle q,u\rangle$ is $\mu$-strongly convex and $L$-smooth on the bounded domain where $\tilde{p}(u)\in[\varepsilon,1-\varepsilon]^{P}$. For step size $\eta\in(0,1/L]$, gradient or mirror descent satisfies
$\mathcal{L}(u_{t+1})-\mathcal{L}(u^{\star})\;\leq\;(1-\mu\eta)\big(\mathcal{L}(u_{t})-\mathcal{L}(u^{\star})\big),$ (18)
yielding linear convergence within each fixed-anchor phase.
Proof sketch.
The Hessian $\nabla^{2}A(u)=\operatorname{Diag}(\tilde{p})-\tilde{p}\,\tilde{p}^{\top}$ has eigenvalues in $\min_{i}\tilde{p}_{i}(1-\tilde{p}_{i}),\max_{i}\tilde{p}_{i}(1-\tilde{p}_{i})]$. On the bounded domain $\varepsilon,1-\varepsilon]^{P}$, this yields:
  • Strong convexity: $\mu\geq\varepsilon(1-\varepsilon)$
  • Smoothness: $L\leq 1/4$
Standard convergence rates for strongly convex and smooth functions then give the stated linear convergence [28]. ∎
Corollary 3.11 (Convergence rate).
With optimal step size $\eta=1/L$, the number of iterations to reach $\varepsilon$-accuracy is:
$T=O\!\Big(\frac{L}{\mu}\log\frac{1}{\varepsilon}\Big)\;=\;O\!\Big(\frac{1}{\varepsilon(1-\varepsilon)}\log\frac{1}{\varepsilon}\Big).$
For well-separated distributions (large $\varepsilon$), convergence is fast.
Dynamic anchor: Coordinate frame changes.
When the anchor updates ($\pi_{\mathrm{ref}}\!\leftarrow\!\pi_{\theta}$), the coordinate frame changes and monotonicity is not guaranteed across updates. However, local stability is preserved.
Theorem 3.12 (Local stability under dynamic anchor).
Let $\pi_{\mathrm{ref}}^{(t)}$ denote the anchor at iteration $t$, and define anchored logits $u^{(t)}=(\ell_{\theta}-\ell_{\mathrm{ref}}^{(t)})/\tau_{\mathrm{anc}}$. Within each anchor phase $t_{k},t_{k+1})$, the loss $\mathcal{L}^{(k)}(u^{(t)})$ decreases monotonically (Theorem 3.10).
At anchor update $t_{k+1}$, the coordinate system shifts, and the loss $\mathcal{L}^{(k+1)}(u^{(k+1)})$ is measured in the new coordinate frame. Global monotonicity $\mathcal{L}^{(k+1)}\leq\mathcal{L}^{(k)}$ is not guaranteed, but the Fisher-metric constraint ensures local stability: the policy remains within a trust region around the previous policy.
Remark 3.13 (Comparison to TRPO/PPO).
TRPO [23] enforces monotonic improvement through explicit KL constraints, while PPO [4] uses clipping to approximate this. Online ADPO achieves similar stability through implicit Fisher geometry, without requiring explicit constraints or clipping. The key difference is that ADPO operates in distribution space (anchored coordinates), while TRPO/PPO operate in parameter space.
3.7 Algorithm and Implementation
We now provide a unified algorithmic description of ADPO that covers both online RL and offline preference learning through a mode parameter.
Algorithm: ADPO (Unified).
Algorithm 1 ADPO (Unified): Online RL and Offline Preference Learning
1:Mode $\in\{\textsc{online},\textsc{offline}\}$; temperature $\tau_{\mathrm{anc}}$; learning rate(s) $\eta_{\pi}$ (and $\eta_{Q}$ for online); anchor update period $N_{\mathrm{anchor}}$; reward temperature $\beta_{r}$ (online only)
2:Initialize policy $\pi_{\theta}$; set anchor $\pi_{\mathrm{ref}}\leftarrow\pi_{\theta}$
3:if online then
4:  Initialize critic $Q_{\phi}$, replay buffer $\mathcal{B}$
5:else
6:  Load dataset $\mathcal{D}=\{(x,S_{x},\text{supervision})\}$
7:end if
8:for training loop $t=1,\ldots,T$ $\triangleright$ iteration for online; epoch for offline do
9:  if online then
10:   Collect trajectories using $\pi_{\theta}$, add to $\mathcal{B}$
11:  end if
12:  for batch update do
13:   if online then
14:     Sample states $\{x_{j}\}_{j=1}^{B}\sim\mathcal{B}$
15:   else
16:     Sample batch $\{(x_{j},S_{x_{j}},\text{sup}_{j})\}_{j=1}^{B}\sim\mathcal{D}$
17:   end if
18:   for each context $x_{j}$ in batch do
19:     // Build target distribution $q$
20:     if online then
21:      Sample candidates $\{y_{1},\ldots,y_{P}\}$: $y_{i}\sim\pi_{\theta}(\cdot|x_{j})$ or mixed sampling
22:      Compute scores: $R_{i}\leftarrow Q_{\phi}(x_{j},y_{i})$ or $r(x_{j},y_{i})$
23:      Normalize (Eq. 19): $\tilde{R}_{i}\leftarrow(R_{i}-\mathrm{median}(R))/(\mathrm{MAD}(R)+\varepsilon)$
24:      Build target: $q(i|x_{j})\leftarrow\mathrm{softmax}(\tilde{R}/\beta_{r})_{i}$
25:     else
26:      Use candidate set $S_{x_{j}}$ from data
27:      Compute target: $q(i|S_{x_{j}})$ from Bradley-Terry, Plackett-Luce, reward model, etc.
28:     end if
29:     $\triangleright$ ===== Core ADPO Update =====
30:     Compute log-probabilities: $\ell_{i}\leftarrow\log\pi_{\theta}(y_{i}|x_{j})$, $\ell_{i}^{\mathrm{ref}}\leftarrow\log\pi_{\mathrm{ref}}(y_{i}|x_{j})$
31:     $u_{i}\leftarrow(\ell_{i}-\ell_{i}^{\mathrm{ref}})/\tau_{\mathrm{anc}}$
32:     $q$-centering (Eq. 10): $\bar{u}_{i}\leftarrow u_{i}-\sum_{k\in S_{x_{j}}}q(k|x_{j})\,u_{k}$
33:     $\tilde{p}_{\theta}(i|x_{j})\leftarrow\mathrm{softmax}(\bar{u})_{i}$
34:   end for
35:   Compute loss: $\mathcal{L}_{\pi}\leftarrow-\frac{1}{B}\sum_{j=1}^{B}\sum_{i\in S_{x_{j}}}q(i|x_{j})\log\tilde{p}_{\theta}(i|x_{j})$
36:   Update policy: $\theta\leftarrow\theta-\eta_{\pi}\nabla_{\theta}\mathcal{L}_{\pi}$
37:  end for
38:  if online then
39:   Update critic $Q_{\phi}$ using standard RL (e.g., TD, SAC)
40:  end if
41:  // Update anchor
42:  if $t\bmod N_{\mathrm{anchor}}=0$ then
43:   $\pi_{\mathrm{ref}}\leftarrow\pi_{\theta}$ $\triangleright$ dynamic anchor (TRPO-style)
44:  end if
45:end for
46:return $\pi_{\theta}$
Practical notes.
  • $q$-centering (always): Apply Eq. 10 to center anchored logits at target $q$, ensuring numerical stability and Fisher-metric consistency.
  • Reward normalization (online RL only): In online mode (Algorithm 1), standardize rewards before building $q$:
    $\tilde{R}_{i}=\frac{R_{i}-\mathrm{median}(R)}{\mathrm{MAD}(R)+\varepsilon},\qquad q(i|x)=\mathrm{softmax}(\tilde{R}/\beta_{r}),$ (19)
    where $\mathrm{MAD}(R)=\mathrm{median}(|R_{i}-\mathrm{median}(R)|)$. Not needed in offline mode where $q$ is fixed.
  • Anchor caching: Cache $\ell^{\mathrm{ref}}$ on batch data to avoid redundant forward passes.
  • Numerical stability: Use log-sum-exp trick for $\log\sum_{j}\exp(u_{j})$.
  • Temperature: In KD scenarios, $\tau_{\mathrm{anc}}\equiv T$ (teacher temperature); see Table 1.
  • Batch processing: Process groups independently; no cross-group dependencies.
4 Experimental Setup
Evaluation metrics.
We use three primary metrics to assess preference learning quality:
  • WinMass: For each context $x$ with candidate set $S_{x}=\{1,\ldots,P\}$ and ground-truth ranking (higher reward is better), WinMass measures the total probability mass assigned to the top item under the anchored distribution: $\mathrm{WinMass}=\mathbb{E}_{x}[\tilde{p}_{\theta}(\operatorname*{arg\,max}_{i}R_{i}^{\star}\mid S_{x})]$. Higher is better; random baseline = $1/P$. Computed over 1000 test contexts per scenario.
  • NDCG@k: Normalized Discounted Cumulative Gain at position $k$, a standard ranking metric that rewards placing high-reward items at top positions with logarithmic discounting. Higher is better; range $[0,1]$. We report NDCG@4 for $P=4$ candidates.
  • KL divergence: For distillation tasks, we measure $\mathrm{KL}(\pi_{\mathrm{teacher}}\|\pi_{\mathrm{student}})$, the expected KL divergence from teacher to student policy over states and actions. Lower is better, indicating closer alignment with the teacher. Computed over 10K states sampled from the replay buffer.
Reporting standards.
All results report mean $\pm$ standard error over 10 random seeds (5 seeds for continuous control). We verify statistical significance using Wilcoxon signed-rank tests (p<0.05) for all claimed improvements; detailed confidence intervals and p-values are provided in the Appendix. Training steps and sample counts are aligned across all methods within each experiment.
Student–teacher KL.
We report the group-averaged divergence
$\mathbb{E}_{x,S_{x}}\Big[\mathrm{KL}\big(\pi_{T}(\cdot\mid S_{x})\|\pi_{\theta}(\cdot\mid S_{x})\big)\Big],$
where both policies are renormalized on the candidate set $S_{x}$ before computing KL.
$2\times 2$ base design + listwise extensions. We systematically compare:
  • Anchoring: Standard DPO (no anchoring) vs. ADPO (anchored to reference policy)
  • Label type: Soft ($q_{ij}\in(0,1)$ via Bradley–Terry) vs. Hard (winner=1, loser=0)
The $2\times 2$ base covers pairwise methods (4 combinations), plus 2 ADPO listwise extensions (Raw/KDE aggregating full distributions), yielding 6 methods total: Standard DPO Pairwise-Soft/Hard, ADPO Pairwise-Soft/Hard, ADPO Listwise-Raw/KDE.
Scenarios and difficulty levels. We test 4 noise types $\times$ 3 severity levels:
  1. (i)
    Heavy Noise: Gaussian noise with outliers. Light (noise scale=1.0, 5% outliers), Medium (scale=2.0, 10%), Heavy (scale=2.5, 20%).
  2. (ii)
    Distribution Shift: Train/test distribution mismatch. Light (scale=1.2, shift=0.3), Medium (1.5, 0.5), Heavy (2.0, 1.0).
  3. (iii)
    Adversarial: Maliciously flipped labels. Light (5%), Medium (10%), Heavy (20%).
  4. (iv)
    Heavy-Tailed: Cauchy noise. Light (scale=0.3), Medium (0.5), Heavy (1.0).
Scenario generation details.
For each prompt $x$ with context $c$ and items $\{v_{i}\}$, rewards are $R_{i}^{\star}=f_{\star}(c,v_{i})$ (MLP). We corrupt observations $\tilde{R}_{i}$ as follows. Heavy Noise: $\tilde{R}_{i}=R_{i}^{\star}+\epsilon_{i}$, $\epsilon_{i}\sim\mathcal{N}(0,\sigma^{2})$ with $p_{\mathrm{out}}$ i.i.d. outliers from $\mathcal{N}(0,\sigma_{\mathrm{out}}^{2})$. Distribution Shift: train uses $(c,v_{i})$; test uses $(\alpha c+\delta,\,v_{i})$ with $\alpha>1,\ \delta\neq 0$. Adversarial: with rate $p$, flip pairwise winners when forming labels/soft targets. Heavy-Tailed: $\epsilon_{i}\sim\mathrm{Cauchy}(0,\gamma)$. All tables report test WinMass under the shifted/noisy process, with $P=4$ fixed throughout.
Model architecture. Policy is an MLP: $\ell_{i}=\text{MLP}(\text{concat}(c,v_{i}))$ where $c\in\mathbb{R}^{D_{c}}$ is context, $v_{i}\in\mathbb{R}^{D_{v}}$ is item embedding. We test 3 scales:
  • Small: hidden=64, layers=2 (total $\sim$8K params)
  • Medium: hidden=128, layers=3 (total $\sim$50K params)
  • Large: hidden=256, layers=4 (total $\sim$260K params)
Candidate set size. All experiments use $P=4$ candidates per group unless otherwise noted.
Training. We train for 80 epochs with batch size 32, learning rate $5\times 10^{-4}$ (no decay), AdamW optimizer ($\beta_{1}=0.9$, $\beta_{2}=0.999$, weight decay $10^{-4}$). Reference policy pre-trained for 30 steps on clean data using the same optimizer configuration; both Standard DPO and ADPO initialize the trainable policy from identical random seeds to ensure fair comparison.
Metrics. WinMass: expected probability mass on the true-best item, i.e., $\mathbb{E}[\tilde{p}_{\theta}(i^{*}|S)]$ where $i^{*}$ is the optimal item. Random baseline = $1/P=0.25$ for $P=4$. All results report mean $\pm$ std over 10 random seeds.
5 Results and Analysis
5.1 $2\times 2$ Base Comparison Across Scenarios
Across all 12 scenarios, ADPO shows consistent relative gains over the standard DPO baseline, with improvements ranging from 12% to 93% (10 seeds). The magnitude of gains increases with noise severity, and listwise training attains the highest end-state performance in 11/12 settings. Detailed per-scenario statistics, confidence intervals, and significance tests are provided in the Appendix.
Table 2: WinMass across 12 scenarios (random baseline = 1/P = 0.25 for P=4). Values are mean over 10 random seeds (std << 0.05 for most entries; 95% CI and Wilcoxon p-values in Appendix). Bold: best method. Underline: best among pairwise. "Best Listwise" shows the highest-performing variant among ADPO Listwise-Raw/KDE, with superscript: $^{\text{R}}$=Raw, $^{\text{K}}$=KDE. In our controlled synthetic noise settings, ADPO shows relative improvements ranging from 12% to 93% over Standard DPO baseline (mean of Std-Soft/Hard).
Scenario Difficulty Std-Soft Std-Hard ADPO-Soft ADPO-Hard Best Listwise
[…]
Figure 2: Comprehensive $2\times 2$ comparison across 12 scenarios (10 seeds each). Each subplot shows convergence curves for 6 methods. Key findings: (i) ADPO methods (solid orange/red/green lines) consistently outperform Standard DPO (solid/dashed blue lines) across all scenarios; (ii) listwise methods achieve highest final performance in 11/12 scenarios; (iii) performance gap widens as difficulty increases (left to right within each row), reaching 93% improvement under Heavy Noise-Heavy; (iv) anchored methods show faster convergence and higher stability. Error bands: mean $\pm$ s.e.
Figure 2 visualizes convergence across all scenarios, revealing consistent ADPO dominance.
5.2 Listwise vs. Pairwise Methods
Listwise methods (ADPO Listwise-Raw/KDE) achieve the highest WinMass in 11 out of 12 scenarios (Table 2). Peak performance: 0.826 (Distribution Shift-Medium, Listwise-Raw). Listwise methods:
  • Use full group information (all $P$ items) vs. pairwise’s $O(P)$ sampled pairs.
  • Benefit from reference anchoring’s groupwise shift invariance (Lemma 2.1).
  • Achieve higher final performance but sometimes converge slower (see Figure 2).
Soft vs. hard labels: context-dependent trade-offs.
Table 3: Soft vs. hard label comparison (pairwise methods only). Winner highlighted. Hard labels dominate under heavy noise (7/12), while soft labels excel under distribution shift and moderate scenarios.
Scenario Difficulty Std-Soft Std-Hard ADPO-Soft ADPO-Hard
Heavy Noise Heavy 0.375 0.418 0.725 0.675
Dist. Shift Light 0.653 0.717 0.752 0.738
Adversarial Medium 0.636 0.643 0.730 0.744
Heavy-Tailed Heavy 0.513 0.508 0.722 0.714
Hard wins: 7/12 scenarios (Heavy Noise 2/3, Adversarial 2/3, Heavy-Tailed 2/3)
Soft wins: 5/12 scenarios (Dist. Shift 2/3, Heavy Noise 1/3, others)
The soft vs. hard label comparison reveals a nuanced picture (Table 3). Hard labels win in 7/12 scenarios, particularly under Heavy Noise (2/3), Adversarial (2/3), and Heavy-Tailed (2/3) settings. However, soft labels excel under distribution shift (2/3 scenarios) and achieve the single highest improvement: ADPO-Soft reaches 0.725 vs. ADPO-Hard’s 0.675 under Heavy Noise-Heavy (+7.4%). Interpretation: The choice depends on noise type—hard labels provide decisive signals when noise is extreme and random, while soft labels preserve gradient information beneficial for systematic distribution shifts and moderate noise.
5.3 Model Scaling
Figure 3: Model scale comparison (Heavy Noise-Medium, 10 seeds). ADPO’s advantage grows with model capacity. Small model: +23% (0.516 vs. 0.420). Medium: +62% (0.716 vs. 0.440). Large: +73% (0.718 vs. 0.416). Standard DPO degrades slightly with scale (overfitting noisy labels), while ADPO benefits from capacity through anchoring. Error bands: mean $\pm$ s.e.
Figure 3 shows larger models amplify ADPO’s benefits. At hidden=256, ADPO-Pairwise-Soft achieves 0.718 vs. Standard DPO’s 0.416 (73% relative gain). Key observation: Standard DPO degrades with scale (Small: 0.420 $\to$ Medium: 0.440 $\to$ Large: 0.416), indicating overfitting risk under noisy labels—larger capacity memorizes corrupted patterns. In contrast, ADPO benefits from increased capacity (Small: 0.516 $\to$ Large: 0.718), confirming that anchoring acts as an effective trust-region regularizer (Lemma 3.6): the Fisher metric $\mathrm{Diag}(q)-qq^{\top}$ constrains policy updates around the reference, preventing overfitting while enabling beneficial capacity utilization.
5.4 Dynamic vs Fixed Anchor Update Strategies
A central design question is whether the reference anchor should be fixed (pre-trained and frozen) or dynamic (periodically updated to match the current policy, similar to TRPO’s old policy). We hypothesize that these strategies serve different purposes: dynamic anchors suit online exploration (where the policy evolves while learning from noisy data), while fixed anchors suit offline distillation (where the policy learns from a stable teacher). Our controlled experiments on online learning scenarios provide a systematic empirical comparison.
Experimental setup.
We compare two anchor update strategies in online exploration scenarios where the policy learns from noisy preference data:
  • ADPO Fixed: Reference $\pi_{\text{ref}}$ pre-trained for 30 steps on light noise, then frozen throughout the 80-epoch training—maintaining a stable geometric center.
  • ADPO Moving: Reference updated every 5 epochs via hard update $\pi_{\text{ref}}\leftarrow\pi_{\theta}$—tracking policy evolution similar to TRPO’s old-policy mechanism.
Both methods use identical model architecture (Medium, 128-dimensional hidden layers), listwise loss formulation, and evaluation protocol, isolating the effect of the anchor update strategy in the online exploration context.
Table 4: Fixed vs Moving Anchor in online exploration scenarios (WinMass). Dynamic anchoring outperforms fixed in 3 out of 4 noise-heavy scenarios where policy evolution is beneficial, achieving 5–11% improvements. Fixed performs slightly better under distribution shift where stability helps generalization. Random baseline = 0.25 for $P=4$ candidates.
Scenario ADPO Fixed ADPO Moving Difference
Heavy Noise (Gaussian + outliers) 0.702 0.780 +11.1%
Distribution Shift (train/test mismatch) 0.706 0.698 -1.1%
Adversarial (label flips) 0.740 0.785 +6.1%
Heavy-Tailed (Cauchy noise) 0.709 0.759 +7.0%
Winner 1/4 3/4
Results and analysis.
Table 4 shows that in online exploration scenarios, dynamic anchoring outperforms fixed in 3 out of 4 cases: heavy noise (+11.1%), adversarial (+6.1%), and heavy-tailed (+7.0%). Only under distribution shift does fixed anchoring perform slightly better (+1.1%).
Why does dynamic anchoring suit online exploration? In noisy environments, the policy evolves as it learns to filter noise and identify robust patterns. A dynamic anchor that tracks this evolution provides an adaptive trust region, allowing the policy to explore more effectively while maintaining stability through the Fisher-metric constraint. This connects ADPO to TRPO [23], which also benefits from tracking policy evolution via old-policy anchors.
However, under distribution shift, a fixed anchor performs slightly better, suggesting that stable geometric centers can aid generalization. More critically, fixed anchors serve a different purpose: as we show in Section 5.5, fixed anchoring excels at offline distillation where the goal is to faithfully learn from a stable teacher, achieving best returns on both continuous control environments with dramatically lower KL divergence. This reveals complementary strengths:
  • Dynamic anchor $\to$ online exploration (adapting to noisy, evolving environments)
  • Fixed anchor $\to$ offline distillation (stable targets for teacher learning)
Ablation: Anchoring mechanism itself.
To isolate the value of anchoring from the choice of update strategy, we compare ADPO (with anchoring) vs ADPO No Anchor (listwise without anchoring). Results show anchoring provides substantial gains across all scenarios: Heavy Noise (+51.2%), Distribution Shift (+3.5%), Adversarial (+10.1%), Heavy-Tailed (+24.0%). This confirms that the anchoring mechanism itself is crucial, with the update strategy determining its best application context.
5.5 Anchoring to Teacher Policies: Continuous Control Distillation
To validate ADPO’s anchor flexibility and test the hypothesis that fixed anchors excel at offline distillation, we conduct experiments on continuous control policy distillation using MuJoCo environments. We evaluate on HalfCheetah-v5 and Hopper-v5; Walker2d-v5 is excluded as the teacher achieved poor performance ($18.86\pm 5.75$), making it unsuitable as a distillation baseline. Here, the student learns from a pre-trained teacher via a frozen (fixed) anchor, contrasting with the online exploration scenarios of Section 5.4 where dynamic anchors proved superior.
Experimental setup.
We train teacher policies using Soft Actor-Critic (SAC) with large networks (512-512-256 hidden dimensions) until convergence. For each environment, we collect a distillation dataset of 10K states, each associated with $P=8$ action candidates sampled from: (i) teacher policy with varying noise levels (75%); (ii) random exploration (25%). Actions are ranked using the teacher’s Q-function, providing preference labels without expensive rollouts. Students are smaller networks (256-256 hidden dimensions) trained for 200 epochs with batch size 128.
We compare three methods:
  • Knowledge Distillation (KD): Standard behavioral cloning minimizing $\mathrm{KL}(\pi_{\mathrm{teacher}}\|\pi_{\theta})$—direct probability matching without anchoring.
  • ADPO-Self-Anchor: Student anchored to its own random initialization $\pi_{\mathrm{init}}$, learning from ranked action preferences.
  • ADPO-Self-Anchor-EMA: Student anchored to an exponential moving average (EMA) of its own policy, providing a slowly moving reference that tightens coupling while remaining stable.
Table 5: Policy distillation on continuous control (MuJoCo). Results averaged over 5 seeds per environment. ADPO methods achieve superior returns while maintaining dramatically lower KL divergence. ADPO-self-anchor uses fixed initial policy as anchor; ADPO-self-anchor-EMA uses exponential moving average. Note: Anchor strategy shows task-dependent effectiveness—fixed anchor excels on HalfCheetah, moving anchor on Hopper.
Environment Method Return NDCG KL $\downarrow$
HalfCheetah-v5 (Teacher: 8476.5) KD $309.0\pm 98.2$ 0.857 30.50
ADPO-self-anchor $\mathbf{279.3\pm 53.2}$ 0.765 10.45
ADPO-self-anchor-EMA $166.8\pm 227.0$ 0.858 0.62
Hopper-v5 (Teacher: 1169.8) KD $36.9\pm 20.2$ 0.799 16.98
ADPO-self-anchor $177.3\pm 158.4$ 0.684 7.61
ADPO-self-anchor-EMA $\mathbf{855.6\pm 269.1}$ 0.755 3.97
Results and analysis.
Table 5 validates our hypothesis that anchoring excels at offline distillation, with task-dependent anchor strategies:
  1. 1.
    ADPO dramatically outperforms KD on both environments. On HalfCheetah-v5, ADPO-self-anchor achieves 279.3 vs KD’s 309.0-309.0 (588-point improvement), while on Hopper-v5, ADPO-self-anchor-EMA reaches 855.6 (73% of teacher’s 1169.8) vs KD’s 36.9 (3%). This demonstrates that ranking quality does not equal decision-making quality—KD achieves high NDCG (0.857, 0.799) but fails to learn effective policies.
  2. 2.
[…]
  3. 3.
    Task-dependent anchor strategy confirmed. Fixed anchor (ADPO-self-anchor) achieves best return on HalfCheetah (279.3), while moving anchor (ADPO-self-anchor-EMA) excels on Hopper (855.6). Both achieve lower KL than KD, with moving anchor providing the tightest coupling (0.62 and 3.97). This validates the theoretical prediction from Section 3.5.
  4. 4.
    Anchor strategy trade-off validated. These distillation results complement the online exploration results of Section 5.4, confirming our hypothesis:
    • Online exploration $\to$ dynamic anchor adapts to policy evolution
    • Offline distillation $\to$ both fixed and moving anchors excel, with task-dependent preferences
Connection to theory.
These results directly validate our geometric framework (Section 3.2, Lemma 3.6):
  • KD vs. ADPO: KD operates in absolute probability space without a fixed geometric center. ADPO anchors the coordinate system at $\pi_{\mathrm{ref}}$, yielding a Fisher-metric trust region.
  • Anchor flexibility: Whether anchoring to an old policy (online RL) or to a fixed reference (e.g., the initial policy or an EMA of the student in offline distillation), the Fisher metric $\operatorname{Diag}(q)-qq^{\top}$ automatically constrains updates. This explains why ADPO succeeds in both settings—the Riemannian geometry is invariant to anchor choice.
  • Preference learning: By optimizing log-odds ratios $\log(\pi_{\theta}/\pi_{\mathrm{ref}})$ to match preference probabilities, ADPO preserves relative action quality rather than absolute probabilities. This is more robust to teacher imperfections and enables students to surpass teachers when beneficial.
Practical guidance.
Table 6: Method selection guide based on empirical results.
Scenario Recommended Method Expected Gain
Heavy Noise ADPO Pairwise-Hard +62–79%
Distribution Shift ADPO Pairwise-Soft +16% (light), listwise +14% (medium)
Adversarial ADPO Listwise-Raw +20–38%
Heavy-Tailed ADPO Listwise-Raw/KDE +30–74%
General (unknown noise) ADPO Listwise-Raw Robust across all scenarios
Knowledge Distillation ADPO-Self-Anchor Superior returns & stability (e.g., 855.6 vs 36.9 on Hopper; 279.3 vs -309.0 on HalfCheetah)
6 Limitations and Future Work
Controlled experimental settings.
Our evaluation uses synthetic contextual bandits with controlled noise processes and continuous control tasks with MuJoCo (8K–260K parameters). Although such setups enable clear causal attributions (e.g., between anchoring and label softness), they may not fully capture real-world annotator heterogeneity or semantic ambiguity. Extrapolation to LLM-scale RLHF (billions of parameters, natural language, human feedback) requires validation on real datasets. We partly mitigate this via multiple noise families (Gaussian with outliers, heavy-tailed, adversarial flips, distribution shift) and multi-seed reporting with significance tests, yet external validity remains an open question.
Limited noise models.
We test Gaussian, adversarial flips, and Cauchy noise. Real-world data may have structured noise (systematic annotator biases, semantic ambiguity, temporal drift). Future work should evaluate ADPO on human preference datasets with genuine uncertainty.
Computational cost.
Listwise methods require computing scores for all $P$ items vs. pairwise’s $K$ samples. For large $P$ (e.g., $P=100$ in ranking), this becomes expensive. Investigating efficient approximations (e.g., top-$k$ subsampling, importance weighting) is a promising direction.
Fairness of reference pre-training.
ADPO methods use a reference policy pre-trained for 30 steps on clean data. This may give ADPO an unfair advantage. However, ablation shows ADPO retains 8–15% gains even with random reference (no pre-training), suggesting anchoring provides value beyond initialization quality. The anchoring structure, not the reference quality, is the key factor.
Future directions for theory and practice.
  • Group size and sampling: We fix $P=4$. Scaling to $P\in\{4,8,16,32\}$ could test: (i) whether larger $P$ amplifies listwise advantage; (ii) whether uncertainty-weighted sampling $\propto q_{ij}(1-q_{ij})$ improves pairwise efficiency.
  • Reference update strategies: We use frozen references. Comparing frozen vs. EMA slow-update ($\tau_{\text{ema}}=0.99$) and varying pre-train steps $N\in\{0,10,30,100\}$ could further isolate "anchoring mechanism" from "better initialization."
  • LLM-scale validation: While we demonstrate ADPO’s effectiveness on contextual bandits and continuous control, validation on large-language model RLHF with human feedback remains an important open question. Our geometric framework suggests ADPO should benefit from iterative refinement (anchoring to previous best policy) and continual learning (anchoring to domain-specific priors).
  • Multi-anchor and dynamic anchoring: Our theory shows anchor flexibility—future work could explore multi-anchor ensembles or adaptive anchor selection during training.
7 Conclusion
We introduced Anchored Direct Preference Optimization (ADPO), a unified projection framework that reveals fundamental connections across major learning paradigms. By formulating policy learning as an anchored projection—minimizing $\mathrm{KL}(q\|\tilde{p}_{\theta})$ with anchored logits $(\ell-\ell^{\mathrm{ref}})/\tau_{\mathrm{anc}}$—we showed that supervised fine-tuning, knowledge distillation, maximum-entropy RL policy updates, advantage-weighted methods, and DPO all emerge as special cases through different choices of target distribution, anchor policy, and temperature. This unification demonstrates that anchoring is not merely a technique for preference learning, but a general principle for robust policy optimization.
[…]
These findings provide clear, actionable guidance: for online RLHF, use dynamic anchors; for offline knowledge transfer, use fixed anchors; for maximum-entropy RL, use uniform anchors; for advantage-weighted updates, use old-policy anchors. By establishing ADPO as a unifying framework with principled anchor selection, we provide a versatile tool for diverse learning scenarios—from noisy preference optimization to stable knowledge distillation. While we have validated our framework on controlled bandits and continuous control, applying this unified anchoring perspective to large-language model alignment, where different training stages (SFT, RLHF, continual learning) may benefit from different anchor strategies, is a critical and promising direction for future work.
Appendix A Connection to Offline RL Methods (AWR/AWAC/MPO)
ADPO connects to advantage-weighted reinforcement learning methods at the level of anchored distributions, though not at the policy level. We detail this partial equivalence here.
A.1 Advantage-Weighted Projection in ADPO
When $\pi_{\mathrm{ref}}=\pi_{\text{old}}$ (old policy anchor), $q(a|s)\propto\pi_{\text{old}}(a|s)\exp\!\big(A^{\pi_{\text{old}}}(s,a)/\beta\big)$ where $A^{\pi_{\text{old}}}=Q^{\pi_{\text{old}}}-V^{\pi_{\text{old}}}$, and $\tau=\beta$, ADPO’s anchored distribution matches:
$\tilde{p}_{\theta}(a|s)=\mathrm{softmax}\!\left(\frac{\log\pi_{\theta}(a|s)-\log\pi_{\text{old}}(a|s)}{\beta}\right)\;\propto\;\pi_{\text{old}}(a|s)\exp\!\big(A/\beta\big).$
Derivation of the induced policy: By Proposition 3.2, the optimal $\pi_{\theta}$ satisfies $\pi_{\theta}\propto\pi_{\mathrm{ref}}\cdot q^{\tau}$. Substituting:
$\pi_{\theta}(a|s)$ $\propto\pi_{\text{old}}(a|s)\cdot\big[\pi_{\text{old}}(a|s)\exp(A/\beta)\big]^{\beta}$
$=\pi_{\text{old}}(a|s)\cdot\pi_{\text{old}}(a|s)^{\beta}\cdot\exp(A)$
$=\pi_{\text{old}}(a|s)^{1+\beta}\cdot\exp\!\big(A^{\pi_{\text{old}}}(s,a)\big).$
A.2 Comparison with Classic AWR/AWAC/MPO
Classic AWR/AWAC M-step [24, 25]: Given advantage-weighted samples, the policy update solves
$\max_{\theta}\mathbb{E}_{(s,a)\sim d}\!\big[\exp(A(s,a)/\beta)\log\pi_{\theta}(a|s)\big],$
which (ignoring normalization) yields $\pi_{\theta}(a|s)\propto\exp(A(s,a)/\beta)$ when capacity is sufficient.
MPO E/M-step [26]: The E-step forms $q(a|s)\propto\exp(Q(s,a)/\eta)$ with dual variable $\eta$ (enforcing a KL budget via Lagrangian). The M-step minimizes $\mathrm{KL}(q\|\pi_{\theta})$, yielding $\pi_{\theta}=q$ under sufficient capacity.
A.3 Key Distinction
ADPO’s equivalence is at the $\tilde{p}-$level (anchored distribution):
$\tilde{p}_{\theta}\;\propto\;\pi_{\text{old}}\exp(A/\beta)\qquad\text{(matches AWR target)},$
but the induced policy $\pi_{\theta}$ differs:
$\text{ADPO:}\quad\pi_{\theta}\;\propto\;\pi_{\text{old}}^{1+\beta}\exp(A),\qquad\text{AWR/MPO:}\quad\pi_{\theta}\;\propto\;\pi_{\text{old}}\exp(A/\beta)\text{ or }\pi_{\theta}=q.$
Interpretation: ADPO’s anchored projection induces a multiplicative fusion $\pi_{\mathrm{ref}}\cdot q^{\tau}$, which preserves the geometric structure of anchoring but differs from classic policy-space updates. The power $1+\beta$ arises from the composition: anchoring at $\pi_{\text{old}}$ followed by temperature-scaled projection toward $q$.
When does this matter? In tabular settings or with full expressivity over discrete actions, both formulations can represent the same optimal behavior (via different parametrizations). In function approximation with shared parameters, the difference becomes significant: ADPO’s form provides an implicit trust region via Fisher geometry, while AWR relies on explicit advantage weighting.
Appendix B Fisher Metric Connection and Relation to KD/TRPO
Fisher metric expansion and Bregman divergence perspective.
For logits $u_{i}=(s_{i}-s_{i}^{\mathrm{ref}})/\tau$, consider the anchored KL objective
$\mathrm{KL}(q\;\|\;\tilde{p}_{\theta})\quad\text{with}\quad\tilde{p}_{\theta}(i)=\frac{\exp(u_{i})}{\sum_{j}\exp(u_{j})}.$
Define the convex function $F(u)=A(u)-\langle q,u\rangle$ where $A(u)=\log\sum_{j}e^{u_{j}}$ is the log-partition function. Then the anchored cross-entropy can be written as a Bregman divergence:
$\mathrm{KL}(q\;\|\;\tilde{p}_{\theta})\;=\;F(u)-F(u^{\star})-\langle\nabla F(u^{\star}),u-u^{\star}\rangle,$ (20)
where $u^{\star}$ satisfies $\tilde{p}_{\theta}(\cdot;u^{\star})=q$ (equivalently, $\nabla F(u^{\star})=0$). The Bregman divergence (20) is always non-negative, induces a three-point identity (Pythagorean theorem for Bregman divergences), and provides a unified view of mirror descent and anchored projection updates.
Expanding around $u^{\star}$, the second-order Taylor expansion yields
$\mathrm{KL}(q\;\|\;\tilde{p}_{\theta})~=~\tfrac{1}{2}\,\delta^{\top}\!\big(\operatorname{Diag}(q)-qq^{\top}\big)\,\delta~+~o(\|\delta\|^{2}),\qquad\delta=u-u^{\star}.$ (21)
The matrix $\operatorname{Diag}(q)-qq^{\top}$ is exactly the Fisher information of the softmax family (the Hessian $\nabla^{2}A(u^{\star})$), revealing that the anchored KL objective induces a local quadratic form corresponding to the Fisher metric in the distribution space. Therefore, ADPO provides an implicit trust region whose shape is governed by this Fisher geometry, with connections to mirror descent and Bregman proximal methods.
Relation to KD and TRPO.
Table 7 summarizes the geometric distinctions. While knowledge distillation (KD) minimizes $\mathrm{KL}(q\|p_{\theta})$ to directly align absolute probabilities, it lacks any notion of a trust region and is sensitive to teacher noise. Trust-region policy optimization (TRPO) instead constrains $\mathrm{KL}(\pi_{\theta_{\text{old}}}\|\pi_{\theta})$ in the parameter space, explicitly bounding the step size at the cost of solving a constrained subproblem. ADPO can be interpreted as a distribution-space analogue of TRPO: by anchoring logits to a reference policy, its anchored KL naturally induces the same Fisher information geometry without requiring an explicit KL penalty. This yields a ‘soft” trust region regularization directly in the output space.
Table 7: Geometric comparison among KD, TRPO, and ADPO.
Method Optimization space Geometric center Fisher metric Trust region
KD Probability simplex Teacher $q$ Softmax Fisher None
TRPO Parameter space Old policy Parameter Fisher Explicit constraint
ADPO Distribution (anchored) Reference policy Softmax Fisher Implicit (anchored KL)
Appendix C Proofs and Derivations
C.1 Proof of Proposition 3.4
We prove that ADPO unifies major learning paradigms through different choices of target $q$, anchor $\pi_{\mathrm{ref}}$, and temperature $\tau$.
(i) Supervised fine-tuning (SFT).
When $\pi_{\mathrm{ref}}$ is uniform (constant logits $s_{i}^{\mathrm{ref}}=c$ for all $i$), the anchored distribution becomes
$\tilde{p}_{\theta}(i\mid S_{x})=\frac{\exp((s_{i}-c)/\tau)}{\sum_{j}\exp((s_{j}-c)/\tau)}=\frac{\exp(s_{i}/\tau)}{\sum_{j}\exp(s_{j}/\tau)}=\mathrm{softmax}_{i}(s_{i}/\tau).$
If $q$ is the data label distribution (one-hot or soft) and $\tau=1$, then $\tilde{p}_{\theta}=\pi_{\theta}$ and $\mathcal{L}_{\mathrm{ADPO}}=\mathbb{E}[\mathrm{CE}(q,\pi_{\theta})]$, the standard cross-entropy loss. For $\tau\neq 1$, the temperature acts as a scaling parameter on the logits (temperature scaling).
(ii) Knowledge distillation (KD).
If $q=\pi_{T}$ (teacher) and $\pi_{\mathrm{ref}}$ is uniform, then with a training temperature $T>1$ we set $\tau=T$ and minimize
$\mathcal{L}_{\mathrm{ADPO}}=\mathbb{E}\!\big[\mathrm{KL}\big(\pi_{T}\;\|\;\mathrm{softmax}(s_{\theta}/T)\big)\big],$
and multiply the loss by $T^{2}$ to recover the classic distillation objective $T^{2}\,\mathrm{CE}\!\big(\pi_{T},\mathrm{softmax}(s_{\theta}/T)\big)$ [22]. Note that $\mathrm{KL}\!\big(\pi_{T}\|\mathrm{softmax}(s_{\theta}/T)\big)=\mathrm{KL}(\pi_{T}\|\pi_{\theta})$ only when $T=1$. If instead $\pi_{\mathrm{ref}}=\pi_{T}$ (anchored KD), we obtain
$\mathcal{L}_{\mathrm{ADPO}}=\mathbb{E}\!\big[\mathrm{KL}\big(\pi_{T}\;\|\;\mathrm{softmax}\!\big((s_{\theta}-s_{T})/\tau\big)\big)\big],$
which performs optimization in teacher-anchored coordinates with an implicit Fisher-metric trust region (see Lemma 3.6).
(iii) Maximum-entropy RL.
For $q(a\mid s)\propto\exp(Q(s,a)/\alpha)$ and uniform $\pi_{\mathrm{ref}}$, minimizing $\mathcal{L}_{\mathrm{ADPO}}$ yields
$\tilde{p}_{\theta}(a\mid s)=\mathrm{softmax}_{a}(s_{a}/\tau)\propto\exp(Q(s,a)/\alpha),$
so $s_{a}/\tau\approx Q(s,a)/\alpha+\mathrm{const}$. Choosing $\tau=1$ gives $s_{a}\approx Q(s,a)/\alpha$, hence $\pi_{\theta}\propto\exp(Q/\alpha)$, the soft-optimal Boltzmann form.
(iv) DPO (binary comparison form).
For $|S_{x}|=2$ with candidates $\{i,j\}$ and binary $q(i)=1,q(j)=0$ (or Bradley–Terry soft targets $q(i)=\sigma(\beta_{r}\Delta R)$), the ADPO loss becomes
$\mathcal{L}_{\mathrm{ADPO}}=-q(i)\log\tilde{p}_{\theta}(i)-q(j)\log\tilde{p}_{\theta}(j),$
where $\tilde{p}_{\theta}(i)=\sigma((\Delta_{\theta}-\Delta_{\mathrm{ref}})/\tau)$ with $\Delta_{\theta}=s_{i}-s_{j}$ and $\Delta_{\mathrm{ref}}=s_{i}^{\mathrm{ref}}-s_{j}^{\mathrm{ref}}$. For uniform $\pi_{\mathrm{ref}}$ (so $\Delta_{\mathrm{ref}}=0$) and hard labels, this reduces to
$\mathcal{L}_{\mathrm{ADPO}}=-\log\sigma(\Delta_{\theta}/\tau),$
which matches the original DPO objective with inverse temperature $1/\tau$. ∎
C.2 Proof of Lemma 3.6
Setup.
Let $u_{i}=(s_{i}-s_{i}^{\mathrm{ref}})/\tau$ and $\tilde{p}_{\theta}(i;u)=\exp(u_{i})/\sum_{j}\exp(u_{j})$. Define $A(u)=\log\sum_{j}e^{u_{j}}$ and $\mathcal{L}(u)=A(u)-\langle q,u\rangle$. Note $\nabla A(u)=\tilde{p}_{\theta}(\cdot;u)$ and $\nabla^{2}A(u)=\mathrm{Diag}(\tilde{p}_{\theta})-\tilde{p}_{\theta}\tilde{p}_{\theta}^{\top}$. Also $\mathcal{L}(u)=\mathrm{KL}\big(q\|\tilde{p}_{\theta}(\cdot;u)\big)+H(q)$.
Second-order expansion.
Let $u^{\star}$ satisfy $\tilde{p}_{\theta}(\cdot;u^{\star})=q$. Then $\nabla\mathcal{L}(u^{\star})=0$ and
$\mathcal{L}(u)=\mathcal{L}(u^{\star})+\tfrac{1}{2}(u-u^{\star})^{\top}\!\big[\mathrm{Diag}(q)-qq^{\top}\big](u-u^{\star})+o\!\big(\|u-u^{\star}\|^{2}\big).$
Centering.
Softmax is invariant to additive shifts: $\tilde{p}_{\theta}(\cdot;u+c\mathbf{1})=\tilde{p}_{\theta}(\cdot;u)$. Hence $\mathrm{Diag}(q)-qq^{\top}$ has a null-space along $\mathbf{1}$. Define $q$-centered logits $\delta_{i}=u_{i}-\sum_{j}q_{j}u_{j}$. Then
$\mathcal{L}(u)-\mathcal{L}(u^{\star})=\tfrac{1}{2}\,\delta^{\top}\!\big[\mathrm{Diag}(q)-qq^{\top}\big]\delta+o(\|\delta\|^{2})=\tfrac{1}{2}\,\operatorname{Var}_{i\sim q}[\delta_{i}]+o(\|\delta\|^{2}).$
Back to anchored scores.
Since $u=(s-s^{\mathrm{ref}})/\tau$, we obtain
$\mathrm{KL}\big(q\|\tilde{p}_{\theta}\big)\approx\frac{1}{2\tau^{2}}\,\operatorname{Var}_{i\sim q}\!\big[(s_{i}-s_{i}^{\mathrm{ref}})-\mathbb{E}_{q}[s-s^{\mathrm{ref}}]\big],$
i.e., a local quadratic trust region in the softmax Fisher metric, with radius scaled by $1/\tau^{2}$. ∎
C.3 Proof of Lemma 2.1
Anchored student invariance.
For any constant $c\in\mathbb{R}$, replace $s_{i}\mapsto s_{i}+c$ and $s_{i}^{\mathrm{ref}}\mapsto s_{i}^{\mathrm{ref}}+c$ for all $i$ in a group $S_{x}$. Then $(s_{i}-s_{i}^{\mathrm{ref}})$ is unchanged, hence
$\tilde{p}_{\theta}(i|S_{x})=\frac{\exp\!\big((s_{i}-s_{i}^{\mathrm{ref}})/\tau\big)}{\sum_{j\in S_{x}}\exp\!\big((s_{j}-s_{j}^{\mathrm{ref}})/\tau\big)}$
is invariant.
Plackett–Luce teacher invariance.
Let $q(i|S_{x})\propto\exp(\hat{R}_{i}/\beta_{r})$. For any constant $c\in\mathbb{R}$, if $\hat{R}_{i}\mapsto\hat{R}_{i}+c$ for all $i\in S_{x}$, then both numerator and denominator are multiplied by the same factor $e^{c/\beta_{r}}$, so $q(i|S_{x})$ is unchanged.
Loss invariance.
Since both $q(\cdot|S_{x})$ and $\tilde{p}_{\theta}(\cdot|S_{x})$ are invariant to groupwise additive shifts, the anchored listwise cross-entropy $\sum_{i}q(i|S_{x})\log\tilde{p}_{\theta}(i|S_{x})$ is invariant as well. ∎
Appendix D Implementation and Training Details
D.1 Noise Generation Procedures
We provide complete mathematical specifications for the four noise families used in our experiments.
Heavy Noise (Gaussian with outliers).
For each item $i$ in group $S_{x}$, the observed noisy reward is:
$\tilde{R}_{i}=R_{i}^{\star}+\epsilon_{i},$
where $R_{i}^{\star}=f_{\star}(c,v_{i})$ is the ground-truth reward from the oracle MLP $f_{\star}$, and
$\epsilon_{i}\sim\begin{cases}\mathcal{N}(0,\sigma^{2})&\text{with probability }1-p_{\text{out}},\\ \mathcal{N}(0,\sigma_{\text{out}}^{2})&\text{with probability }p_{\text{out}},\end{cases}$
with $\sigma_{\text{out}}^{2}=100\sigma^{2}$. The severity levels correspond to:
  • Light: noise scale $\sigma=1.0$, $p_{\text{out}}=0.05$
  • Medium: noise scale $\sigma=2.0$, $p_{\text{out}}=0.10$
  • Heavy: noise scale $\sigma=2.5$, $p_{\text{out}}=0.20$
Distribution Shift.
Training and test distributions differ in the context space. Let $(c,v_{i})$ denote training instances. Test instances are generated as:
$(c_{\text{test}},v_{i})=(\alpha\cdot c+\delta\cdot\mathbf{1},v_{i}),$
where $\alpha>1$ scales the context magnitude and $\delta$ adds a constant shift. Severity levels:
  • Light: $\alpha=1.2$, $\delta=0.3$
  • Medium: $\alpha=1.5$, $\delta=0.5$
  • Heavy: $\alpha=2.0$, $\delta=1.0$
The ground-truth model $f_{\star}$ trained on the original distribution must generalize to these shifted contexts.
Adversarial Label Flips.
For pairwise comparisons $(i,j)$, we flip the preference label with probability $p$:
$y_{ij}^{\text{flip}}=\begin{cases}1-y_{ij}&\text{with probability }p,\\ y_{ij}&\text{with probability }1-p,\end{cases}$
where $y_{ij}=\mathbb{I}(R_{i}^{\star}>R_{j}^{\star})$ is the clean label. For soft labels, we apply:
$q_{ij}^{\text{flip}}=\begin{cases}1-q_{ij}&\text{with probability }p,\\ q_{ij}&\text{with probability }1-p.\end{cases}$
Severity levels: Light (5%), Medium (10%), Heavy (20%).
Heavy-Tailed Noise (Cauchy).
The observed reward is:
$\tilde{R}_{i}=R_{i}^{\star}+\epsilon_{i},\quad\epsilon_{i}\sim\text{Cauchy}(0,\gamma),$
where $\text{Cauchy}(0,\gamma)$ has PDF $f(\epsilon)=\frac{1}{\pi\gamma(1+(\epsilon/\gamma)^{2})}$. This distribution has undefined mean and variance, making it particularly challenging. Severity levels:
  • Light: $\gamma=0.3$
  • Medium: $\gamma=0.5$
  • Heavy: $\gamma=1.0$
D.2 Hyperparameters and Training Configuration
Table 8: Complete hyperparameter configuration for all experiments.
Parameter Value
Optimization
Optimizer AdamW
Learning rate $5\times 10^{-4}$
$\beta_{1}$ 0.9
$\beta_{2}$ 0.999
Weight decay $10^{-4}$
Learning rate schedule Constant (no decay)
Batch size 32
Training epochs 80
Model Architecture
Context dimension $D_{c}$ 8
Item embedding $D_{v}$ 8
Small model hidden 64
Medium model hidden 128
Large model hidden 256
Small model layers 2
Medium model layers 3
Large model layers 4
Activation ReLU
ADPO-specific
Temperature $\beta$ 1.0
Temperature $\beta_{r}$ 1.0
Listwise temperature $\tau$ 1.0 (pairwise), grid-search (listwise)
Reference pre-train steps 30 (on clean data)
Experimental (Contextual Bandit)
Candidate set size $P$ 4
Number of seeds 10
Training dataset size 10,000 instances
Test dataset size 1,000 instances
Distillation Experiment (MuJoCo)
Teacher network SAC with 512-512-256 hidden dims
Student network 256-256 hidden dims
Action candidates $K$ 8
Distillation dataset 10,000 states
Training epochs 200
Batch size 128
Number of seeds 5
Teacher sampling ratio 75% (teacher) + 25% (random)
D.3 Metrics Definition
WinMass.
WinMass is defined as the expected probability mass assigned to the ground-truth best item:
$\mathrm{WinMass}=\mathbb{E}_{x\sim\mathcal{D}_{\text{test}},S_{x}}\left[\tilde{p}_{\theta}(i^{\star}|S_{x})\right],$
where $i^{\star}=\operatorname*{arg\,max}_{i\in S_{x}}R_{i}^{\star}$ is the item with the highest ground-truth reward. For random selection with $P=4$ candidates, the baseline WinMass is $1/P=0.25$.
Appendix E Additional Experiments and Ablations
E.1 Statistical Significance Tests
Method.
For each scenario and method, we report mean WinMass over 10 random seeds. We compute 95% confidence intervals using the bootstrap method with 10,000 resamples. To test whether ADPO significantly outperforms Standard DPO, we use the paired Wilcoxon signed-rank test on the 10-seed results. The null hypothesis is that the median difference is zero. We report $p$-values and reject $H_{0}$ at $\alpha=0.05$.
E.2 Effect of Reference Initialization
Ablation results.
We evaluate ADPO performance with varying reference pre-training steps $N\in\{0,10,30,100\}$ on Heavy Noise-Medium scenario. Results show ADPO maintains 8–15% advantage even with random reference ($N=0$), confirming that the anchoring structure itself matters beyond initialization quality. Full pre-training ($N=100$) provides marginal additional gains (+2–3%), suggesting that geometric centering is the key factor rather than reference quality alone.
E.3 Temperature Sensitivity (Full Grid)
Temperature sensitivity analysis shows that ADPO maintains robust performance across a wide range of temperature values. Key findings:
  • Pairwise-ADPO maintains WinMass $\in[0.61,0.62]$ across all $(\beta_{r},\tau)\in\{0.5,1,2,4\}^{2}$;
  • Listwise-Raw shows similar robustness (WinMass $\in[0.58,0.63]$);
  • Listwise-KDE benefits from lower $\tau$ (0.5–1.0) but degrades at $\tau>2.0$.
E.4 KDE-CDF-Logit Transform Details
Given rewards $\{\tilde{R}_{i}\}_{i=1}^{P}$ in a group, ADPO-KDE constructs targets via:
  1. 1.
    KDE fit (Scott’s rule). Fit a Gaussian KDE $\hat{f}(r)=\frac{1}{Ph}\sum_{i=1}^{P}K\!\big(\frac{r-\tilde{R}_{i}}{h}\big)$ with $K$ standard normal and bandwidth $h=P^{-1/5}\,\hat{\sigma}$ (Scott’s rule), where $\hat{\sigma}$ is the sample standard deviation.
  2. 2.
    CDF evaluation. Compute $\hat{F}(\tilde{R}_{i})=\int_{-\infty}^{\tilde{R}_{i}}\hat{f}(r)\,dr$.
  3. 3.
    Logit map with clipping. Let $\hat{F}_{\epsilon}(\tilde{R}_{i})=\mathrm{clip}\big(\hat{F}(\tilde{R}_{i}),\,10^{-6},\,1-10^{-6}\big)$ and $\ell_{i}=\log\!\frac{\hat{F}_{\epsilon}(\tilde{R}_{i})}{1-\hat{F}_{\epsilon}(\tilde{R}_{i})}$.
  4. 4.
    Temperature-softmax target. Center logits via $\bar{\ell}=\frac{1}{P}\sum_{j}\ell_{j}$ and set
    $q(i|S_{x})=\frac{\exp\!\big((\ell_{i}-\bar{\ell})/\beta_{r}\big)}{\sum_{j}\exp\!\big((\ell_{j}-\bar{\ell})/\beta_{r}\big)}.$
This rank-like transform bounds outlier influence through the CDF and preserves separability via the logit.
Appendix F Reproducibility Checklist
To facilitate reproduction of our results, we provide the following information:
Random seeds.
All experiments use seeds $\{0,1,\ldots,9\}$ for 10-seed runs. Each seed controls:
  • PyTorch random number generator: torch.manual_seed(seed)
  • NumPy random number generator: np.random.seed(seed)
  • Python built-in random: random.seed(seed)
Dataset generation.
Synthetic contextual bandits are generated as follows:
  • Context $c\sim\mathcal{U}[0,1]^{D_{c}}$ with $D_{c}=8$;
  • Item embeddings $v_{i}\sim\mathcal{U}[0,1]^{D_{v}}$ with $D_{v}=8$;
  • Ground-truth rewards $R_{i}^{\star}=f_{\star}(\text{concat}(c,v_{i}))$ where $f_{\star}$ is a 2-layer MLP with hidden dimension 64;
  • Noisy observations $\tilde{R}_{i}$ generated according to Section D.
Fixed random generation ensures identical train/test splits across all methods.
Code and configuration.
Code and configuration files are provided in the supplementary materials. The repository includes:
  • Implementations of all ADPO variants (pairwise/listwise, soft/hard);
  • Standard DPO baseline implementation;
  • Noise generation scripts for all four families;
  • Experimental configuration YAMLs;
  • Plotting scripts for all figures;
  • Pre-computed results for verification.
Evaluation procedure.
For each trained model:
  1. 1.
    Generate 1,000 test instances using the same distribution as training (or shifted distribution for Distribution Shift scenarios);
  2. 2.
    For each instance, sample $P=4$ candidates;
  3. 3.
    Compute WinMass as $\mathbb{E}[\tilde{p}_{\theta}(i^{\star}|S_{x})]$ over all test instances;
  4. 4.
    Report mean and standard deviation over 10 random seeds.